### v2

# Gaussian approximation on the Skorokhod space via Malliavin calculus and regularization

Abstract.
We introduce a carré du champ operator for Banach-valued random elements, taking values in the projective tensor product, and use it to control the bounded Lipschitz distance between a Malliavin-smooth random element satisfying mild regularity assumptions and a Radon Gaussian taking values in the Skorokhod space equipped with the uniform topology. In the case where the random element is a Banach-valued multiple integral, the carré du champ expression is further bounded by norms of the contracted integral kernel. The main technical tool is an integration by parts formula, which might be of independent interest. As a by-product, we recover a bound obtained recently by Düker and Zoubouloglou in the Hilbert space setting and complement it by providing contraction bounds.
Key words and phrases:
Banach spaces; Malliavin calculus; $\gamma$-radonifying operators; Projective tensor product; Multiple Wiener–Itô integrals; Skorokhod space; Bounded Wasserstein distance; Functional Gaussian approximation.
2010 Mathematics Subject Classification:
60G15, 60H07, 60F17, 60B11, 62E17
S. Bourguin is supported by the Simons Foundation (Grant 635136).
1. Introduction
Gaussian quantitative approximation via Malliavin integration by parts is a well-established technique which led to significant developments in several branches in mathematics and has found numerous applications. In a nutshell, starting from a so called integral probability distance, i.e. a metric $d$ on the class of probability measures of the form
$d(P_{F},P_{Z})=\sup\left|\mathbb{E}(g(F))-\mathbb{E}(g(Z))\right|,$
where $F$ is a sufficiently smooth Malliavin differentiable random vector, $Z$ a multivariate Gaussian (both centered for simplicity), the supremum is taken over a suitable class of test functions, one transforms the right hand side via Stein’s method, the so-called smart path method or another suitable technique into a functional expression in $F$, and then uses integration by parts to arrive at a bound of the form
(1) $d(F,Z)\leq c\,\left\lVert\Gamma(F,-L^{-1}F)-R\right\rVert,$
where the matrix $\Gamma(F,-L^{-1}F)=(\Gamma(F_{i},-L^{-1}F_{j}))_{1\leq i\leq j\leq n}$ is given by $\Gamma(F_{i},-L^{-1}F_{j})=\left\langle DF_{i},-DL^{-1}Fj\right\rangle_{\mathfrak{H}}$, with $D$ the Malliavin derivative, $L^{-1}$ the pseudo-inverse of the generator and $\mathfrak{H}$ the Hilbert space where the underlying isonormal Gaussian process is defined. The pioneering and pathbreaking work [NP09] of Nourdin and Peccati in dimension one was generalized to multiple but finite dimension by the same authors and Reveillac shortly afterwards (see [NPR10]). For a book-length treatment of the finite-dimensional theory, we refer to  [NP12], its far reaching ramifications in diverse areas of mathematics are well-known and classical.
In the context of functional approximation, the finite-dimensional theory has successfully been used to provide convergence of finite-dimensional distributions, while tightness had to be taken care of by other means. The resulting limit theorems are therefore typically non-quantitative. In order to obtain quantitative results from the method in this setting, an analogue of the bound (1) in infinite-dimension is required. The main obstacle when trying to generalize the known proofs for the finite-dimensional theory is an exploding dimension dependent constant $c$.
In [BC20], the authors of the present article proposed an analogue of (1) with a dimension-free constant for random elements taking values in a separable Hilbert space, controlling a smooth distance by the norm of the difference of the carré du champ $\Gamma$ and the covariance operator $R$, interpreted as random square integrable Hilbert-Schmidt operators. However, as was noticed recently (see  [BBCP25]), the particular smooth distance used does not metrize convergence in distribution, severely limiting the use of the resulting bounds and invalidating the proofs of all weak convergence claims made in  [BC20]. It directly follows from the discussion in [BBCP25] that the only reasonable parameter choice (yielding a bound on a distance which does metrize convergence in distribution on separable Hilbert spaces) leads to the trace instead of the Hilbert-Schmidt norm. This observation has already been leveraged in [DZ25] to prove a non-quantitative Fourth Moment Theorem for multiple Wiener-Itô integrals taking values in separable Hilbert spaces.
In the present article, we illuminate the infinite-dimensional situation further by introducing a Banach-valued carré du champ operator $\Gamma$ associated to a Malliavin-smooth random element $X$, and use it to control the bounded Lipschitz distance on the Skorokhod space (or any closed subspace thereof) equipped with the uniform norm between a Radon Gaussian measure and the random element $X$. Our starting point is the Banach-valued Malliavin calculus introduced by Maas and van Neerven (see  [MvN08], [Maa10], [PV14]). The role of the trace norm is taken over by its natural generalization, which is the projective norm on the projective tensor product of the Banach space with itself. As is well known, if the Banach space has the approximation property, the projective tensors are isometrically isomorphic to the nuclear operators, which in turn coincide with the trace class operators in the Hilbert space case.
As is known from the general theory of probability in Banach spaces, all random elements taking values in a Banach space $E$ and having a strong second moment possess nuclear covariance operators $R_{\gamma}\colon E^{\ast}\to E$, which can be seen as elements of the projective tensor product $E\widehat{\otimes}_{\pi}E$. The carré du champ operator $\Gamma$ associated to a Malliavin-smooth random element $X$ takes values in the same tensor product, and, as in finite dimension, its projection on the constants (the cokernel of the expectation operator) coincides with the covariance operator of $X$. If $X$ is Gaussian, then $\Gamma$ is deterministic.
Heuristically, as in the finite-dimensional case, one should hence continue to think of $\Gamma$ as a random covariance operator, and the distance-controlling quantity $\left\lVert\Gamma-R_{\gamma}\right\rVert_{L^{2}(\Omega;E\widehat{\otimes}_{\pi}E)}$ as measuring its randomness. This heuristic becomes particularly lucid when investigating the special case where $X$ is an $E$-valued multiple Wiener-Itô integral as constructed in [Maa10]. We prove that then - like in the finite-dimensional case - $\Gamma$ expands into a sum of the covariance operator of $X$ and multiple integrals whose kernels are now vector-valued contractions of the original kernels of $X$ and take values in $E\widehat{\otimes}_{\pi}E$. Replacing $X$ by a sequence $X_{n}$, in order for weak convergence to a Gaussian to take place, these multiple integrals should vanish in the limit. In finite-dimension, quantifying this heuristic by controlling the distance between $X$ and $Z$ in a suitable metric by the norms of the kernel contractions (which by the isometry property of multiple Wiener-Itô integrals are directly related to their asymptotic behaviour) is arguably one of the most powerful features of the integration by parts approach. We provide an infinite-dimensional analogue of such a contraction bound through use of the vector-valued analogue of the Itô-isometry proved in [Maa10].
To explain our results in more details, we prove an abstract carré du champ bound of the form
(2) $\left|\mathbb{E}(f(X))-\mathbb{E}(f(Z))\right|\leq\frac{1}{2}\left\lVert f^{\prime\prime}\right\rVert_{\infty}\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{\pi},$
where $f$ is twice Fréchet differentiable, $X$ is a Malliavin-differentiable random element, $Z$ a Radon Gaussian, both taking values in a Banach space $E$, and $Q_{Z}$ denotes the covariance tensor of $Z$. The norm on the right is the projective norm. See Proposition 3.10 for a rigorous statement. Once the carré du champ has been constructed, the remaining ingredients in order to obtain (2) are an integration by parts formula, which might be of independent interest, which can then be coupled with the so-called smart-path method.
At this point, it should be mentioned that the class of Fréchet differentiable functions is not rich enough to metrize convergence in distribution on general Banach spaces. However, confining ourselves to the Skorokhod space $\mathcal{D}$ equipped with the uniform topology and using a regularization procedure developed in  [BRZ24] (which builds upon Barbour’s idea of a pre-limiting Gaussian process; see  [Bar90]), we can leverage the bound (2) to yield a bound for the bounded Lipschitz distance $d_{\operatorname{BL}}$ of the form
(3) $d_{\operatorname{BL}}(X,Z)\leq\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}\right)+\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)+C_{\varepsilon}\,\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{\mathcal{D}\widehat{\otimes}_{\pi}\mathcal{D}}\right),$
where the Gaussian $Z$ is assumed to have almost surely continuous sample paths and, for $\varepsilon>0$, $X_{\varepsilon}$ and $Z_{\varepsilon}$ denote the so-called $\varepsilon$-regularizations of $X$ and $Z$, respectively, obtained by a local uniform smoothing procedure (see Section 4 for detailed statements and definitions). Under the assumption that also $X$ has almost surely continuous sample paths, one can dispense of the regularization terms and obtain a bound in terms of the $\Gamma$-expression only. For example, if $X$ and $Z$ admit uniform regularizations with Hölder continuous modulus of continuity of order $\alpha$, we obtain
(4) $d_{\operatorname{BL}}(X,Z)\leq c_{\alpha,T}\mathbb{E}\left(\left\lVert\Gamma_{\pi}\left(X,-L^{-1}X\right)-Q_{Z}\right\rVert_{E\widehat{\otimes}_{\pi}E}\right)^{e_{\alpha}},$
with $e_{\alpha}=\alpha/(\alpha+2/3)$.
Of course, as soon as both $X$ and $Z$ take values in a closed subspace $E$, such as the space of continuous functions, the bounds (3) and (4) remain valid for the bounded Lipschitz distance there. Furthermore, if $X$ is a multiple integral (or a finite sum thereof), we can further bound the right hand side of (4) with the norms of the kernel contractions mentioned above.
As a by-product, when specializing to a Hilbert space, the abstract bound (2) immediately yields the trace class bound of [DZ25] for the smooth distance $\rho_{\infty}$, and we complement their result by providing contraction bounds for Hilbert-valued multiple integrals (see Section 7 for details).
We believe that in order to avoid the regularization procedure and obtain bounds on general Banach spaces, it is necessary to leave the class of Fréchet differentiable functions and instead consider the much richer class of $H$-differentiable functions (where $H$ is the Cameron-Martin space of the Gaussian measure). This topic is currently under investigation.
The paper is organized as follows: in Section 2, we gather the necessary preliminaries needed for our analysis, in particular essentials from the theory of topological tensor products, $\gamma$-radonifying operators and Banach valued Malliavin calculus as well as Banach-valued Wiener chaos. The carré du champ operator, integration by parts formulæ and a rigorous statement of the bound (2) is developed in Section 3. In Section 4 we apply the regularization procedure to control the bounded Lipschitz distance in several scenarios. The contraction bound for the norm of $\Gamma-R_{\gamma}$ is developed in Section 5, followed by an application in Section 6 to a Gaussian-subordinated Hermite-model on the space of continuous functions. We finish with a brief discussion of the Hilbert space case in Section 7.
2. Preliminaries
2.1. Generalities
Let $E$ be a real Banach space. If not mentioned otherwise, all occuring random elements are defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$, and $\mathbb{E}$ denotes mathematical expectation with respect to $\mathbb{P}$.
All vector spaces in this paper are assumed to be real. For a Banach space $E$, let $U_{E}$ be its closed unit ball and $E^{\ast}$ its topological dual. If $F$ is another Banach space, their topological direct sum is denoted by $E\oplus F$ and equipped with the norm $\left\lVert u\oplus v\right\rVert=\left\lVert u\right\rVert_{E}+\left\lVert v\right\rVert_{F}$ to turn it into a Banach space. If a closed subspace $U$ of $E$ is complemented by another closed subspace $V$, then $E$ itself is isomorphic to the topological direct sum $U\oplus V$. Note that in particular, the natural projections onto $U$ and $V$ are both continuous.
For vector spaces $E_{1},E_{2},\dots,E_{k},F$, we denote by $L(E_{1},E_{2},\dots,E_{k};F)$ the space of $k$-linear $F$-valued mappings on $E_{1}\times E_{2}\times\dots\times E_{k}$ and by $\mathcal{F}(E_{1},E_{2},\dots,E_{k};F)$ the subspace of finite-rank mappings. If the vector spaces are normed, then $\mathcal{L}(E_{1},E_{2},\dots,E_{k};F)$ denotes those $k$-linear mappings which are bounded (i.e. continuous with respect to the norm $\left\lVert\varphi\right\rVert=\sup_{x_{j}\in U_{j}}\left\lVert\varphi(x_{1},\dots,x_{k})\right\rVert_{F}$), so that in particular $\mathcal{F}(E_{1},\dots,E_{k};F)\subseteq\mathcal{L}(E_{1},\dots,E_{k};F)$. We write $L_{k}(E,F)=L(\underbrace{E,E,\dots,E}_{\text{$k$ times}};F)$ and $L(E,F)=L_{1}(E,F)$, with analogous shorthands for the respective bounded or finite-rank subspaces.
We write $\mathcal{C}^{k}(E,F)$ for the space of $k$-times continuously Fréchet differentiable functions, and $\mathcal{C}_{b}^{k}(E,F)$ for the subspace of those $f\in C^{k}(E,F)$ such that $f$ and its first $k$ derivatives are bounded. The Fréchet derivative of $f$ is denoted by $f^{\prime}$.
2.2. Topological tensor products
We denote by $E\otimes F$ the algebraic tensor product of $E$ and $F$, by $E\otimes_{\alpha}F$ the corresponding normed space, where $\alpha$ is some norm on $E\otimes F$, and by $E\widehat{\otimes}_{\alpha}F$ its completion. A norm $\alpha$ on the class of normed spaces is called a tensor norm (also called uniform cross norm), if $\alpha$ is reasonable, i.e. $\varepsilon\leq\alpha\leq\pi$, where $\varepsilon$ and $\pi$ are the injective and projective norms, respectively, and the metric mapping property is satisfied: for normed spaces $E_{i},F_{i}$ and bounded operators $T_{i}\in\mathcal{L}(E_{i},F_{i})$, $i=1,2$, the tensor operator $T_{1}\otimes T_{2}\colon E_{1}\otimes_{\alpha}E_{2}\to F_{1}\otimes_{\alpha}F_{2}$ is bounded and satisfies
$\left\lVert T_{1}\otimes T_{2}\right\rVert\leq\left\lVert T_{1}\right\rVert\left\lVert T_{2}\right\rVert,$
where all norms are the usual operator norms. In particular, for simple tensors $x\otimes y\in E\otimes F$, one has $\left\lVert x\otimes y\right\rVert_{\alpha}=\left\lVert x\right\rVert_{E}\left\lVert y\right\rVert_{F}$. The Hilbert cross norm $\sigma$ (obtained from the inner product induced by defining $\left\langle x_{1}\otimes y_{1},x_{2}\otimes y_{2}\right\rangle_{\sigma}=\left\langle x_{1},y_{1}\right\rangle\left\langle x_{2},y_{2}\right\rangle$) is a tensor norm in the above sense as well, albeit of course only on pairs of Hilbert spaces. With the above notation, the algebraic tensor product of two Hilbert spaces $\mathfrak{H}$ and $\mathfrak{K}$ equipped with the tensor norm $\sigma$ is denoted by $\mathfrak{H}\otimes_{\sigma}\mathfrak{K}$, the Hilbert space obtained upon completing by $\mathfrak{H}\widehat{\otimes}_{\sigma}\mathfrak{K}$. As is well known, this Hilbert space is isometrically isomorphic to the Hilbert-Schmidt operators.
The symmetric tensor product of $E$ and $F$ is denoted by $E\odot F$. For a positive integer $p$ and an associative tensor norm $\alpha$ (such as the Hilbert cross norm $\sigma$), the notations $E^{\otimes p}$, $E^{\otimes_{\alpha}p}$ and $E^{\widehat{\otimes}_{\sigma}p}$ as well as analogous versions for the symmetric tensor powers have their obvious meaning.
We will mainly be concerned with the projective tensor product $E\widehat{\otimes}_{\pi}F$, which is obtained from the projective tensor norm $\left\lVert\cdot\right\rVert_{\pi}$, given for $u\in E\otimes F$ by
$\left\lVert u\right\rVert_{\pi}=\inf\left\{\sum_{k=1}^{m}\left\lVert x_{k}\right\rVert_{E}\left\lVert y_{k}\right\rVert_{F}\colon u=\sum_{k=1}^{m}x_{k}\otimes y_{k}\text{ in }E\otimes F\right\},$
where the infimum is taken over all finite representations of $u$ as a sum of simple tensors. Any element $u\in E\widehat{\otimes}_{\pi}F$ has representations of the form $u=\sum_{k=1}^{\infty}x_{k}\otimes y_{k}$ with $x_{k}\in E$ and $y_{k}\in Y$ such that $\sum_{k=1}^{\infty}\left\lVert x_{k}\right\rVert_{E}\left\lVert y_{k}\right\rVert_{F}<\infty$. The dual of the projective tensor product can be identified with the space of bounded bilinear forms $\mathcal{L}(E,F;\mathbb{R})$, where the duality pairing $(\varphi,u)_{\pi}$ between $\varphi\in\mathcal{L}(E,F;\mathbb{R})$ and a tensor $u=\sum_{k=1}^{m}x_{k}\otimes y_{k}\in E\otimes F$ is defined as
(5) $(\varphi,u)_{\pi}=\sum_{k=1}^{m}\varphi(x_{k},y_{k}).$
As a tensor norm $\alpha$ satisfies $\alpha\leq\pi$, the dual of $E\widehat{\otimes}_{\alpha}F$ can be identified with the subspace of those bounded bilinear forms $\mathcal{L}(E,F;\mathbb{R})$ which are also continuous with respect to the dual norm defined by $\alpha$. Furthermore, by the dual representation of a norm, it holds that
(6) $\left|(\varphi,u)_{\alpha}\right|\leq\left\lVert\varphi\right\rVert_{\left(E\widehat{\otimes}_{\alpha}F\right)^{\ast}}\,\left\lVert u\right\rVert_{\alpha}.$
For $u\in E\otimes F$ and $\varphi\in\left(E\widehat{\otimes}_{\alpha}F\right)^{\ast}$, the pairings $(\varphi,u)_{\alpha}$ and $(\varphi,u)_{\pi}$ coincide in value and can be computed via (5).
For the projective norm, the dual representation reads
$\left\lVert u\right\rVert_{\pi}=\sup_{\varphi\in U_{{\mathcal{L}(E,F;\mathbb{R})}}}\left|\left(\varphi,u\right)_{\pi}\right|$
and we can write (6) more explicitly as
$\left|(\varphi,u)_{\pi}\right|\leq\left\lVert\varphi\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}\left\lVert u\right\rVert_{\pi}$
for all $u\in E\widehat{\otimes}_{\pi}F$ and $\varphi\in\mathcal{L}(E,F;\mathbb{R})$.
A bounded linear operator $T\colon E\to F$, with $E$ and $F$ Banach spaces, is called nuclear, if there exist $f_{k}\in E^{\ast}$ and $y_{k}\in F$ such that
$Tx=\sum_{k=1}^{\infty}\left(f_{k},x\right)_{E^{\ast},E}y_{k}$
for all $x\in E$ and
$\sum_{k=1}^{\infty}\left\lVert f_{k}\right\rVert_{E^{\ast}}\,\left\lVert y_{k}\right\rVert_{E}<\infty.$
Such a representation is called nuclear. The nuclear norm $\left\lVert T\right\rVert_{\nu}$ of $T$ is defined as
$\left\lVert T\right\rVert_{\nu}=\inf\sum_{k=1}^{\infty}\left\lVert f_{k}\right\rVert\left\lVert y_{k}\right\rVert,$
where the infimum is taken over all nuclear representations $T=\sum_{k=1}^{\infty}f_{k}\otimes y_{k}$. The space of all nuclear operators is denoted by $\mathfrak{N}(E,F)$. The finite rank operators are densely contained.
By identifying a simple tensor $x\otimes y\in E^{\ast}\otimes F$ with the rank one operator $T_{x\otimes y}\in\mathcal{L}(E,F)$ defined by $T_{f\otimes y}x=f(x)y$, then extending by linearity and continuity, we obtain a metric surjection $J\colon E^{\ast}\widehat{\otimes}_{\pi}F\to\mathfrak{N}(E,F)$. It is straightforward to see, that
(7) $\left\lVert J(u)\right\rVert_{\nu}\leq\left\lVert u\right\rVert_{\pi}.$
Indeed, the infimum in the projective norm is taken over finite representations only, while for the nuclear norm also infinite representations are allowed. As is well known, $J$ might fail to be injective, so that it is in general not possible to identify $E^{\ast}\widehat{\otimes}_{\pi}F$ with $\mathfrak{N}(E,F)$. However, $J$ is always injective on the algebraic tensor product, and for $u\in E^{\ast}\otimes_{\pi}F$ (such that $J(u)$ has finite rank) we have equality in (7).
For the trace on the algebraic tensor product $E^{\ast}\otimes E$ and the finite-rank operators $\mathcal{F}(E,E)$, we will use the symbol $\operatorname{tr}$ (or $\operatorname{tr}_{E}$ if the Banach space should be emphasized). The trace on the projective tensor product $E^{\ast}\widehat{\otimes}_{\pi}E$ will be denoted by $\operatorname{tr}_{\pi}$ or $\operatorname{tr}_{E,\pi}$. If $J$ is injective, $\operatorname{tr}_{\pi}$ induces a trace on $\mathfrak{N}(E)=\mathfrak{N}(E,E)$, for which we will again use $\operatorname{tr}$ or $\operatorname{tr}_{E}$. Recall that the trace on $E^{\ast}\otimes E$ is the unique linearization of the evaluation map on $E^{\ast}\times E$, i.e. the bilinear form defined by $(f^{\ast},x)\mapsto\left(f^{\ast},x\right)_{E^{\ast},E}$. For an algebraic tensor
$u=\sum_{k=1}^{n}f_{k}^{\ast}\otimes x_{k}\in E^{\ast}\otimes E,$
we hence get
$\operatorname{tr}\left(u\right)=\sum_{k=1}^{n}\left(f_{k}^{\ast},x_{k}\right)_{E^{\ast},E}.$
The relevant condition for injectivity of $J$ is the approximation property, which is in particular shared by all Hilbert spaces. If $K$ is a Hilbert space, we can thus identify $K\widehat{\otimes}_{\pi}K$ with $\mathfrak{N}(K)$, and furthermore the nuclear norm coincides with the trace norm in this case, so that $\mathfrak{N}(K)$ is the space of trace class operators. The trace itself can then of course also be computed using an orthonormal basis.
2.3. Banach valued random elements and Bochner spaces
Let $(E,\mathcal{S})$ be a Banach space equipped with a $\sigma$-field $\mathcal{S}$ and $(\Omega,\mathcal{F},\mu)$ a probability space. A random element $X$ on $E$ is a measurable mapping $X\colon\Omega\to E$. Only the cylindrical $\sigma$-field $\mathcal{E}(E)$ (the minimal $\sigma$-field with respect to which all continuous linear functionals on $E$ are measurable) and the Borel $\sigma$-field $\mathcal{B}(E)$ on $E$ will be considered for $\mathcal{S}$ in this article. While these two $\sigma$-fields coincide if $E$ is separable, in general one has $\mathcal{E}(E)\subseteq\mathcal{B}(E)$.
A Borel random element $X$ is called Radon, if its law is Radon, i.e. a Borel probability measure on $E$ which is tight, in the sense that for any $\varepsilon>0$ there exists a compact set $K_{\varepsilon}\subseteq E$ such that $\mu\left(X\in K_{\varepsilon}\right)\geq 1-\varepsilon$. Equivalently, $X$ almost surely takes its values in a closed, separable subspace of $E$. The law of a Radon random element is uniquely determined by its values on $\mathcal{E}(E)$. As usual, we identify two random elements, if they agree almost everywhere.
For $p\in[1,\infty]$, the Bochner spaces of (equivalence classes of) $E$-valued Borel random elements of strong order $p$ will be denoted by $L^{p}(\Omega;E)$ or $L^{p}(\mu;E)$ if we want to emphasize the measure. If $E=\mathbb{R}$, we simply write $L^{p}(\Omega)$ or $L^{p}(\mu)$. As elements of $L^{p}(\Omega;E)$ are almost surely separably valued, they are Radon. Note that $\mathcal{E}(E)$-measurable random elements of strong order $p$ have unique Radon extensions to $\mathcal{B}(E)$ (see  [VTC87, Prop. II.1.1]).
The algebraic tensor product $L^{p}(\Omega)\otimes E$ is embedded in the Bochner space $L^{p}(\Omega;E)$ via $f]_{\mu}\otimes x\mapsto[f\otimes x]_{\mu}$, where $f]_{\mu}$ and $f\otimes x]_{\mu}$ denote the respective equivalence classes. The norm on $L^{p}(\Omega;E)$ thus induces a norm $\Delta_{p}$ on $L^{p}(\Omega)\otimes E$ and it is well known (see for example [DF93, ch. 7]), that $L^{p}(\Omega)\widehat{\otimes}_{\Delta_{p}}E\cong L^{p}(\Omega;E)$, that $\varepsilon\leq\Delta_{p}\leq\pi$ and that $\Delta_{p}$ is not a tensor norm, as it fails the metric mapping property: if $S\in\mathcal{L}(L^{p}(\mu),L^{q}(\nu))$ and $T\in\mathcal{L}(E,F)$, where $\nu$ is another finite measure, $F$ a Banach space and $q\in[1,\infty)$, then the corresponding tensor operator $S\otimes T\colon L^{p}(\mu)\otimes_{\Delta_{p}}E\to L^{q}(\nu)\otimes_{\Delta_{q}}F$ is not necessarily continuous, i.e. it is not guaranteed that $\left\lVert S\otimes T\right\rVert\leq\left\lVert S\right\rVert\,\left\lVert T\right\rVert$.
However, there are several situations where continuity follows (see [DF93, ch. 7] or the survey  [DF92]), one of which will be used in this paper:
Theorem 2.1.
Let $E$, $F$ be Banach spaces, $\mu$, $\nu$ be two measures, $p$, $q\in[1,\infty]$ and $T\in\mathcal{L}(E,F)$. If $S\in\mathcal{L}(L^{p}(\mu),L^{q}(\nu))$ is positive (mapping a.e. positive functions to a.e. positive functions), then $S\otimes T\colon L^{p}(\mu)\otimes_{\Delta_{p}}E\to L^{q}(\nu)\otimes_{\Delta_{q}}F$ is continuous and one has $\left\lVert S\otimes T\right\rVert\leq\left\lVert S\right\rVert\,\left\lVert T\right\rVert$.
A proof can be found in [DF93, p.80].
The cross-covariance operator (see [VTC87, Ch. III.2.4]) $R_{X,Y}\colon F^{\ast}\to E$ of an $E$-valued random element $X$ and an $F$-valued random element $F$, both of strong second order, is defined by
$R_{X,Y}=\mathbb{E}\left((Y-\mathbb{E}(Y),\cdot)_{F,F^{\ast}}\left(X-\mathbb{E}(X)\right)\right)=\mathbb{E}\left((Y,\cdot)_{F,F^{\ast}}\left(X-\mathbb{E}(X)\right)\right).$
As is well known (see for example [VTC87, Ch. III.2.3]), the strong second order of $X$ and $Y$ implies that $R_{X,Y}$ is nuclear.
In the same way, we define the cross covariance tensor $Q_{X,Y}\in F\widehat{\otimes}_{\pi}E$ by
$Q_{X,Y}=\mathbb{E}\left(Y\otimes\left(X-\mathbb{E}(X)\right)\right).$
If $X=Y$, the positive and symmetric covariance operator (resp. tensor) of $X$ is defined by $R_{X}=R_{X,Y}$ (resp. $Q_{X}=Q_{X,X}$).
2.4. Radon Gaussian measures
Let $E$ be a Banach space. A measure $\gamma$ on $\mathcal{E}(E)$ is called Gaussian, if the measure $\gamma\circ f^{-1}$ is Gaussian for any $f\in E^{\ast}$. A Radon measure $\gamma$ on $\mathcal{B}(E)$ is called a Radon Gaussian measure, if its restriction to $\mathcal{E}(E)$ is a Gaussian measure. A Gaussian measure is called centered (resp. (non)-degenerate), if the Gaussian laws induced on the dual elements are centered (resp. (non)-degenerate).
An $E$-valued random element is called (Radon) Gaussian, centered and/or (non)-degenerate, if its induced law has the respective properties. Two Gaussian random elements $X$ and $Y$ taking values in Banach spaces $E$ and $F$, respectively, are jointly Gaussian, if the random element $X\oplus Y\in E\oplus F$, with $E\oplus F$ denoting the Banach direct sum of $E$ and $F$, is Gaussian, i.e. if for all $e^{\ast}\in E^{\ast}$ and $f^{\ast}\in F^{\ast}$, the distribution of $e^{\ast}(X)+f^{\ast}(Y)$ is Gaussian.
Let $\gamma$ be a centered Radon Gaussian measure and let $R_{\gamma}\colon E^{\ast}\to E$ be its covariance operator, defined in analogy to the covariance operator of a random element of strong order two as the Bochner integral
(8) $R_{\gamma}f=\int_{E}f(x)x\gamma(dx).$
Then, one has
$\left\lVert R_{\gamma}\right\rVert_{\nu}\leq\int_{E}\left\lVert x\right\rVert^{2}\gamma(dx).$
Denote by $E_{\gamma}^{\ast}$ the $L^{2}(\gamma)$-completion of the canonical embedding of $E^{\ast}$ into $L^{2}(\gamma)$. As $\left\lVert f\right\rVert_{L^{2}(\gamma)}\leq\int_{E}\left\lVert x\right\rVert_{E}^{2}\gamma(dx)\left\lVert f\right\rVert_{E^{\ast}}$ for all $f\in E^{\ast}$, the operator $R_{\gamma}$ can also be defined (via (8)) on $E^{\ast}_{\gamma}$ and retains its nuclearity. The Cameron-Martin space $H_{\gamma}\subseteq E$ is defined as $R_{\gamma}(E_{\gamma}^{\ast})$. Equipped with the inner product inherited from $L^{2}(\gamma)$, i.e.
$(R_{\gamma}f,R_{\gamma}g)_{H(\gamma)}=(f,g)_{L^{2}(\gamma)}=f\left(R_{\gamma}g\right),$
it becomes a separable Hilbert space which is isometrically isomorphic to $E_{\gamma}^{\ast}$. Moreover, we have the embeddings
$E^{\ast}\subseteq H^{\ast}\simeq H\subseteq E.$
As shown in [KS80] (see [Tar80] for a different proof), one can always find an orthonormal basis $\left\{e_{k}\colon k\in\mathbb{N}\right\}$ of $H_{\gamma}$ such that
$R_{\gamma}=\sum_{k=1}^{\infty}\left(e_{k},\cdot\right)_{E,E^{\ast}}\otimes e_{k}$
is a nuclear representation of $R_{\gamma}$, i.e.
$\sum_{k=1}^{\infty}\left\lVert e_{k}\right\rVert^{2}_{E}<\infty.$
2.5. $\gamma$-radonifying operators
This section and the two following it essentially summarize most of the content of the first five sections of [Maa10], with some additions suitable for our purposes.
An isonormal Gaussian process $W$ on a real separable Hilbert space $\mathfrak{H}$ is a linear isometry $W\colon\mathfrak{H}\to L^{2}(\Omega,P)$, where $(\Omega,\mathcal{F},P)$ is some probability space, such that $W(h)$ is a centered Gaussian random variable for all $h\in\mathfrak{H}$.
The isonormal Gaussian process $W$ induces a Gaussian cylinder set measure on $\mathfrak{H}$, which has no $\sigma$-finite extension unless $\mathfrak{H}$ is finite-dimensional. However, one can always achieve such an extension on a larger Banach space $E$, into which $\mathfrak{H}$ is continuously embedded. Such an embedding is a $\gamma$-radonifying operator.
Definition 2.2.
Let $W$ be an isonormal Gaussian process on a real, separable Hilbert space $\mathfrak{H}$ and let $E$ be a Banach space. A bounded linear operator $T\colon\mathfrak{H}\to E$ is called $\gamma$-radonifying, if there exists a strongly measurable Gaussian random variable $Z\in L^{2}(\Omega;E)$ such that
$W(T^{\ast}f)=f(Z)$
for all $f\in E^{\ast}$, where $T^{\ast}$ denotes the adjoint of $T$ and we have identified $\mathfrak{H}$ with its dual.
A $\gamma$-radonifying operator $T$ as above therefore transports the Gaussian cylindrical distribution on $\mathfrak{H}$ induced by $W$ to a Gaussian cylindrical distribution on $E$, the latter of which can then be extended to a bona fide Radon Gaussian measure $\gamma_{T}$ on $E$.
To connect with the previous section, one can verify directly that the covariance operator $R_{\gamma_{T}}$ is given by $R_{\gamma_{T}}=TT^{\ast}$, where we have identified $\mathfrak{H}$ with its dual. Indeed, for $f,g\in E^{\ast}$ one has
$g(R_{\gamma_{T}}f)=\mathbb{E}\left(f(Z)g(Z)\right)=\mathbb{E}\left(W(T^{\ast}f)W(T^{\ast}g)\right)\\ =\left(T^{\ast}f,T^{\ast}g\right)_{\mathfrak{H}^{\ast}}=\left(TT^{\ast}f,g\right)_{E,E^{\ast}}=g(TT^{\ast}f).$
It can be shown that, given a $\gamma$-radonifying operator $T$, the Gaussian series $\sum_{k=1}^{\infty}W(\mathfrak{h}_{k})T\mathfrak{h}_{k}$ converges in $L^{2}(\Omega;E)$ for some orthonormal basis $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ of $\mathfrak{H}$, and that the convergence and value of the sum does not depend on the choice of the isonormal Gaussian process or the orthonormal basis, so that
$\left\lVert T\right\rVert_{\gamma(\mathfrak{H},E)}=\mathbb{E}\left(\left\lVert\sum_{n}W(\mathfrak{h}_{n})T\mathfrak{h}_{n}\right\rVert_{E}^{2}\right)^{\frac{1}{2}}$
defines a characterizing norm on the space $\gamma(\mathfrak{H},E)$ of all $\gamma$-radonifying operators, turning it into a closed subspace of $\mathcal{L}(\mathfrak{H},E)$. By approximating the above series by its partial sums, it is seen that the finite rank operators $\mathcal{F}(\mathfrak{H},E)$ are densely contained in $\gamma(\mathfrak{H},E)$ and that a $\gamma$-radonifying operator is hence compact. Moreover, $\gamma(\mathfrak{H},E)$ forms an operator ideal, i.e. if $T\in\gamma(\mathfrak{H},E)$, $S\in\mathcal{L}(\mathfrak{G},\mathfrak{H})$ and $R\in\mathcal{L}(E,F)$, where $\mathfrak{G}$ is a Hilbert and $F$ a Banach space, then $RTS\in\gamma(\mathfrak{G},F)$ and $\left\lVert RTS\right\rVert_{\gamma(\mathfrak{G},F)}\leq\left\lVert R\right\rVert_{\mathcal{L}(E,F)}\,\left\lVert T\right\rVert_{\gamma(\mathfrak{H},E)}\,\left\lVert S\right\rVert_{\mathcal{L}(\mathfrak{G},\mathfrak{H})}$.
Inductively, one defines the spaces $\gamma^{p}(\mathfrak{H},E)=\gamma\left(\mathfrak{H},\gamma^{p-1}(\mathfrak{H},E)\right)$ for $p\in\mathbb{N}$, where we set $\gamma^{0}(\mathfrak{H},E)=E$, so that $\gamma^{1}(\mathfrak{H},E)=\gamma(\mathfrak{H},E)$.
The density of the finite rank operators $\mathcal{F}(\mathfrak{H},E)$ in $\gamma(\mathfrak{H},E)$ transfers to the density of $\mathcal{F}^{p}(\mathfrak{H},E)$ in $\gamma(\mathfrak{H},E)$, where $\mathcal{F}^{p}(\mathfrak{H},E)$ is obtained from $E$ analogously to $\gamma^{p}(\mathfrak{H},E)$.
Alternatively, one can view $\gamma(\mathfrak{H},E)$ as the completion of the algebraic tensor product $\mathfrak{H}\otimes_{\gamma}E$ with respect to the tensor norm $\gamma$ (on pairs of Hilbert and Banach spaces) inferred from the identification of $\mathfrak{H}\otimes E$ with $\mathcal{F}(\mathfrak{H},E)$. Then $\gamma(\mathfrak{H},E)\simeq\mathfrak{H}\widehat{\otimes}_{\gamma}E$ and, more generally,
$\gamma^{p}(\mathfrak{H},E)\simeq\mathfrak{H}\,\widehat{\otimes}_{\gamma}\left(\dots\left(\mathfrak{H}\,\widehat{\otimes}_{\gamma}\left(\mathfrak{H}\,\widehat{\otimes}_{\gamma}E\right)\right)\dots\right).$
For simplicity, given $T\in\gamma^{p}(\mathfrak{H},E)$ and $h=(h_{1},\dots,h_{p})\in\mathfrak{H}^{p}$, we write $T(h)=T(h_{1}\dots,h_{p})$ instead of the more cumbersome
$\left(\dots\left(\left(Th_{1}\right)\,h_{2}\right)\dots\right)h_{p}.$
Note that any $T\in\mathcal{F}^{p}(\mathfrak{H},E)$ can be written in the form
(9) $T=\sum_{\mathbf{i}\in[n]^{p}}\mathfrak{h}_{\mathbf{i}}\otimes x_{\mathbf{i}},$
where $n]=\left\{1,2,\dots,n\right\}$, $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ is an orthonormal basis of $\mathfrak{H}$ and, for a multi index $\mathbf{i}=(i_{1},i_{2},\dots,i_{p})\in[n]^{p}$, $x_{\mathbf{i}}\in E$ and the symbol
$\mathfrak{h}_{\mathbf{i}}\otimes x_{\mathbf{i}}$
denotes the rank-one operator in $\mathcal{F}^{p}(\mathfrak{H},E)$ defined for $h=(h_{1},\dots,h_{p})\in\mathfrak{H}^{p}$ by
$\left(\mathfrak{h}_{\mathbf{i}}\otimes x_{\mathbf{i}}\right)h=\prod_{j=1}^{p}\left\langle\mathfrak{h}_{i_{j}},h_{j}\right\rangle_{\mathfrak{H}}x_{\mathbf{i}}.$
The norm for such $T$ is then given by
(10) $\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)}=\mathbb{E}\left(\left\lVert\sum_{\mathbf{i}\in[n]^{p}}W_{\mathbf{i}}\,x_{\mathbf{i}}\right\rVert_{E}^{2}\right)^{\frac{1}{2}},$
where $W\colon\mathfrak{H}\to L^{2}(\Omega)$ is an isonormal Gaussian process and
$W_{\mathbf{i}}=W^{(1)}\left(\mathfrak{h}_{i_{1}}\right)W^{(2)}\left(\mathfrak{h}_{i_{2}}\right)\cdots W^{(p)}\left(\mathfrak{h}_{i_{p}}\right),$
with the $W^{(j)}$ being mutually independent copies of $W$, so that $W_{\mathbf{i}}$ is a product of independent standard Gaussians.
Note that in (10) one can in general not replace the products $W_{\mathbf{i}}$ by $\widetilde{W}(\mathfrak{h}_{\mathbf{i}})$, using a single isonormal Gaussian process $\widetilde{W}$ on $\mathfrak{H}^{\otimes p}$. As shown in  [KW14] (see also  [vNW08]), $\gamma^{p}(\mathfrak{H},E)$ is isomorphic to $\gamma(\mathfrak{H}^{\widehat{\otimes}_{\sigma}p},E)$ for all $p\in\mathbb{N}$ precisely when $E$ has Pisier’s property $(\alpha)$, which was introduced in  [Pis78]. This property is shared in particular by all Hilbert spaces, in which case $\gamma^{p}(\mathfrak{H},E)$ is isometrically isomorphic to the Hilbert tensor product $\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}\widehat{\otimes}_{\sigma}E$ corresponding to multilinear Hilbert-Schmidt mappings.
The symmetrization operator $\operatorname{Sym}\colon\gamma^{p}(\mathfrak{H},E)\to\gamma^{p}(\mathfrak{H},E)$ is defined for $T\in\gamma^{p}(\mathfrak{H},E)$ and $h\in\mathfrak{H}^{p}$ by
$\operatorname{Sym}(T)h=\frac{1}{p!}\sum_{\sigma\in S_{p}}Th_{\sigma},$
where $S_{p}$ is the permutation group on $[p]$ and $h_{\sigma}=\left(h_{\sigma(1)},\dots,h_{\sigma(p)}\right)$. It is clear that $\operatorname{Sym}$ is a projection on $\gamma^{p}(\mathfrak{H},E)$ and we denote its closed range by $\gamma^{p}_{sym}(\mathfrak{H},E)$, calling the operators in this subspace symmetric. Analogously, we define $\mathcal{F}^{p}_{sym}(\mathfrak{H},E)=\operatorname{Sym}\left(\mathcal{F}^{p}(\mathfrak{H},E)\right)$.
For more details on $\gamma$-radonifying operators, see  [vN10] and the references therein.
2.6. Malliavin calculus in Banach spaces
In [Maa10] and [MvN08] (see also  [PV14]), $\gamma$-radonifying operators emerged as natural objects for developing Malliavin calculus for Banach-valued random elements. We will outline the construction of the Malliavin derivative, Ornstein-Uhlenbeck semigroup and its generator from these papers, skipping the divergence operator, which is not needed to prove our results. As we could not find the construction of the pseudo-inverse of the generator in the literature, it is done here in full detail.
We use standard notation for the induced operators of Malliavin calculus, i.e. $D$ (Malliavin derivative), $\delta$ (its adjoint), $(P_{t})_{t\geq 0}$ (Ornstein-Uhlenbeck semigroup), $L$ (its generator) and $L^{-1}$ (the pseudoinverse of the generator) and refer to [Nua06] or [NP12] for details.
For clarity, given a Banach space $E$ different from $\mathbb{R}$, we will indicate throughout this section with a subscript when the Malliavin operators act on $E$-valued random elements, writing $D_{E}$, $L_{E}$ etc. Once all operators have been extended, we will drop the subscript again, which, as explained in the last paragraph of this section, can be done without causing confusion.
Let $W\colon\mathfrak{H}\to L^{2}(\Omega)$ be an isonormal Gaussian process as above and $\mathcal{S}$ be the class of smooth random variables of the form $f(W(h_{1}),\dots,W(h_{n}))$, where $f\in C^{\infty}_{b}(\mathbb{R}^{n})$ is an infinitely differentiable function on $\mathbb{R}^{n}$ such that $f$ and all of its partial derivatives are bounded. For a Banach space $E$, we define $\mathcal{S}(E)=\mathcal{S}\otimes E$ and lift the Malliavin derivative to $\mathcal{S}(E)$ by defining $D_{E}=D\otimes\operatorname{Id}_{E}$. Explicitly, given $F\in\mathcal{S}(E)$ of the form $F=f(W(h_{1}),\dots,W(h_{n}))\otimes x$, with $x\in E$, $D_{E}F$ is the random element
$D_{E}F=\left(Df(W(h_{1}),\dots,W(h_{n}))\right)\otimes x=\sum_{k=1}^{n}\partial_{n}f(W(h_{1}),\dots,W(h_{n}))h_{n}\otimes x,$
where $h_{n}\otimes x$ denotes the rank-one operator $T\colon\mathfrak{H}\to E$ given by $Th=\left\langle h_{n},h\right\rangle_{\mathfrak{H}}x$. This procedure yields closable operators $D_{E}$ mapping from $\mathcal{S}\otimes_{\Delta_{p}}E\subseteq L^{p}(\Omega;E)$ to $\mathcal{S}\otimes_{\Delta_{p}}(\mathfrak{H}\otimes_{\gamma}E)\subseteq L^{p}(\Omega,\gamma(\mathfrak{H},E))$ for any $p\in[1,\infty)$. Setting $D^{1}_{E}=D_{E}$ and, for convenience, $D_{E}^{0}=\operatorname{Id}_{E}$, we can therefore inductively define Malliavin derivatives $D_{E}^{k}\colon L^{p}(\Omega;E)\to L^{p}(\Omega,\gamma^{k}(\mathfrak{H},E))$ for any $k\in\mathbb{N}_{0}$ by $D^{k+1}_{E}F=D_{E}(D^{k}_{E}F)$. The domains of their closures will be denoted by $\mathbb{D}^{k,p}(E)$, together with the corresponding norm
$\left\lVert F\right\rVert_{k,p}=\sum_{j=0}^{k}\left\lVert D^{j}_{E}F\right\rVert_{L^{p}(\Omega;\gamma^{j}(\mathfrak{H},E))},$
and as usual we will continue to use the symbol $D^{k}_{E}$ for the closure. As mentioned in the previous section, if $E$ is a Hilbert space, the space $\gamma^{k}(\mathfrak{H},E)$, coincides with the $k$-multilinear $E$-valued Hilbert-Schmidt mappings. In particular, for $E=\mathbb{R}$ we have $\gamma^{k}(\mathfrak{H},\mathbb{R})\simeq\mathfrak{H}^{\widehat{\otimes}_{\sigma}k}$ and obtain the classical Malliavin derivatives treated for example in [Nua06] or [NP12].
Pronk and Veraar showed in [PV14, Prop. 3.8] that the chain rule has an extension to the vector-valued setting, when the codomain of the smooth transformation is UMD (unconditional for martingale differences). A Banach space $E$ is called UMD, if for some (or equivalently all) $p\in(1,\infty)$ there exists a constant $C_{p}$ such that for all $E$-valued martingale difference sequences $(d_{k})_{k\in\mathbb{N}}$ in $L^{p}(\Omega;E)$ (with respect to some filtered probability space), all binary sequences $(r_{k})_{k\in\mathbb{N}}\subseteq\left\{-1,1\right\}^{\mathbb{N}}$ and all $n\in\mathbb{N}$ one has
$\left\lVert\sum_{k=1}^{n}r_{k}d_{k}\right\rVert_{p}\leq C_{p}\left\lVert\sum_{k=1}^{n}d_{k}\right\rVert_{p}.$
Examples of UMD-spaces are Hilbert spaces and the $L^{p}(\Omega,\mu)$-spaces for $p\in(1,\infty)$ and some $\sigma$-finite measure $\mu$ on a measurable space $(\Omega,\mathcal{F})$. The chain rule reads as follows (see [PV14, Prop. 3.8]).
Theorem 2.3.
Let $E_{1}$ be a Banach space, $E_{2}$ be a UMD Banach space, $p\in(1,\infty)$ and $\varphi\in\mathcal{C}_{b}^{1}(E_{1},E_{2})$. If $F\in\mathbb{D}^{1,p}(E_{1})$, then $\varphi(F)\in\mathbb{D}^{1,p}(E_{2})$ and
$D_{E_{2}}\varphi(F)=\varphi^{\prime}(F)D_{E_{1}}F.$
In the case where $E_{2}=\mathbb{R}$, one may also take $p=1$.
As the positive and strongly continuous Ornstein-Uhlenbeck semigroup $(P_{t})_{t\geq 0}$ is defined on $L^{p}(\Omega)$ for any $p\in[1,\infty)$, the tensor operators
$P_{E,t}=P_{t}\otimes\operatorname{Id}_{E}$
extend to a strongly continuous semigroup $(P_{E,t})_{t\geq 0}$ on the Lebesgue-Bochner spaces $L^{p}(\Omega;E)$ for any Banach space $E$ and $p\in[1,\infty)$ via Proposition 2.4. Let $L_{E}$ be its infinitesimal generator with $L^{p}(\Omega;E)$-domain $\operatorname{dom}_{p}(L_{E})$.
To define $L_{E}^{-1}$, we make use of the fact that the expectation operator (i.e. the Bochner integral) $T_{E}\colon L^{p}(\Omega;E)\to L^{p}(\Omega;E)$ defined by $T_{E}F=\mathbb{E}\left(F\right)=\int_{\Omega}F(\omega)\mu(d\omega)$ is a bounded projection of unit norm for all $p\in[1,\infty)$ and arbitrary Banach spaces $E$. Indeed,
$\left\lVert T_{E}\left(F\right)\right\rVert^{p}_{L^{p}(\Omega;E)}=\mathbb{E}\left(\left\lVert\mathbb{E}\left(F\right)\right\rVert_{E}^{p}\right)\leq\mathbb{E}\left(\left(\mathbb{E}\left(\left\lVert F\right\rVert_{E}\right)^{p}\right)\right)=\left\lVert F\right\rVert^{p}_{L^{p}(\Omega;E)},$
so that $\left\lVert T_{E}\right\rVert_{L^{p}(\Omega;E)}\leq 1$, and evaluating at constants shows that in fact $\left\lVert T_{E}\right\rVert_{L^{p}(\Omega;E)}=1$. As kernel and range of a bounded projection are closed and complement each other, $L^{p}(\Omega;E)\cong\operatorname{ker}(T_{E})\oplus\operatorname{ran}(T_{E})$ is a topological direct sum and the corresponding projections $U_{E}$ and $V_{E}$ onto the kernel and range of $T_{E}$, respectively, are continuous.
On $\operatorname{ker}(T_{\mathbb{R}})\otimes E$, we define $\widetilde{L}^{-1}_{E}F=\int_{0}^{\infty}P_{t,E}Fdt$, where the integral is an improper vector-valued Riemann integral. This integral exists in $L^{p}(\Omega;E)$, as for $F\in\operatorname{ker}(T_{\mathbb{R}})\otimes E$ we have
$\left\lVert P_{t,E}F\right\rVert_{L^{p}(\Omega;E)}=\left\lVert\left(P_{t}\otimes\operatorname{Id}_{E}\right)F\right\rVert_{L^{p}(\Omega;E)}\leq\left\lVert F\right\rVert_{L^{p}(\Omega;E)}\,\left\lVert(P_{t})_{\restriction{\operatorname{ker}(T_{\mathbb{R}})}}\otimes\operatorname{Id}_{E}\right\rVert\\ \leq\left\lVert F\right\rVert_{L^{p}(\Omega;E)}\,\left\lVert(P_{t})_{\restriction{\operatorname{ker}(T_{\mathbb{R}})}}\right\rVert\left\lVert\operatorname{Id}_{E}\right\rVert\leq C_{p}e^{-t}\left\lVert F\right\rVert_{L^{p}(\Omega;E)},$
where $C_{p}$ is a positive constant only depending on $p$ and we have used that $\left\lVert P_{t}G\right\rVert_{L^{p}(\Omega)}\leq C_{p}e^{-t}\left\lVert G\right\rVert_{L^{p}(\Omega)}$ on a dense subset of $\operatorname{ker}(T_{\mathbb{R}})$ (see [Nua06, Lemma 1.4.1]).
As $L_{p}(\Omega)\otimes E$ is dense in $L^{p}(\Omega;E)$ and the projection $U_{E}$ is bounded, $U_{E}(L^{p}(\Omega)\otimes E)=\operatorname{ker}(T_{\mathbb{R}})\otimes E$ is dense in the closure of $U_{E}(L^{p}(\Omega;E))=\operatorname{ker}(T_{E})$. Therefore, $\widetilde{L}_{E}^{-1}$ extends to a bounded operator on $\operatorname{ker}(T_{E})$. Finally, we define $L_{E}^{-1}\colon L^{p}(\Omega;E)\to L^{p}(\Omega;E)$ by $L_{E}^{-1}=\widetilde{L}_{E}^{-1}\circ U_{E}$ and obtain a bounded operator on $L^{p}(\Omega;E)=\operatorname{ker}(T_{E})\oplus\operatorname{ran}(T_{E})$ (as the composition of two bounded operators is bounded).
As the next result shows, the well-known relationship $LL^{-1}F=\mathbb{E}(F)-F$, valid for any $F\in L^{2}(\Omega)$ continues to hold in the vector-valued case, so that $L_{E}^{-1}$ is indeed a pseudo inverse of $L_{E}$.
Proposition 2.4.
For all $F\in L^{p}(\Omega;E)$, $L_{E}^{-1}(F)\in\operatorname{dom}_{p}(L_{E})$ and
$L_{E}L_{E}^{-1}F=\mathbb{E}\left(F\right)-F.$
Proof.
From the general theory of semigroups (see for example  [EN06, Ch. II, Lemma 1.3]), we have for any $F\in L^{p}(\Omega;E)$ and $t>0$ that $\int_{0}^{t}P_{E,s}Fds\in\operatorname{dom}_{p}(L_{E})$ and
$P_{E,t}F-F=L_{E}\int_{0}^{t}P_{E,s}Fds.$
As $P_{E,t}$ maps constants to constants and $L_{E}$ vanishes on constants, we can also write the above identity as
(11) $P_{E,t}F-F=L_{E}\int_{0}^{t}P_{E,s}U_{E}Fds,$
where $U_{E}$ is the projection onto $\operatorname{ker}(T_{E})$ introduced above. As $L_{E}$ is closed and $\int_{0}^{t}P_{E,s}U_{E}Fds\to L_{E}^{-1}F$ as $t\to\infty$, we obtain $L_{E}^{-1}F\in\operatorname{dom}_{p}(L_{E})$ and that the right hand side of (11) converges to $L_{E}L_{E}^{-1}F$ in $L^{p}(\Omega;E)$. It therefore remains to show that $(P_{E,t})_{t\geq 0}$ inherits the asymptotic behaviour of $(P_{t})_{t\geq 0}$, i.e.
$\lim_{t\to\infty}P_{E,t}F=\mathbb{E}(F)$
in $L^{p}(\Omega;E)$. To do so, we approximate $F$ in $L^{p}(\Omega;E)$ by elements $F_{n}\in L^{p}(\Omega)\otimes E$. Then
$\left\lVert P_{E,t}F-\mathbb{E}(F)\right\rVert_{L^{p}(\Omega;E)}$
$\leq\left\lVert P_{E,t}(F-F_{n})\right\rVert_{L^{p}(\Omega;E)}+\left\lVert P_{E,t}F_{n}-\mathbb{E}(F_{n})\right\rVert_{L^{p}(\Omega;E)}+\left\lVert\mathbb{E}(F_{n}-F)\right\rVert_{L^{p}(\Omega;E)}$
(12) $\leq 2\left\lVert F-F_{n}\right\rVert_{L^{p}(\Omega;E)}+\left\lVert P_{E,t}F_{n}-\mathbb{E}(F_{n})\right\rVert_{L^{p}(\Omega;E)},$
where we have used the contraction property of $(P_{E,t})_{t\geq 0}$ and the boundedness of $\mathbb{E}$ to obtain the last inequality. Clearly,
$\lim_{t\to\infty}\left\lVert P_{E,t}F_{n}-\mathbb{E}(F_{n})\right\rVert_{L^{p}(\Omega;E)}=0$
by definition of $P_{E,t}$ and the corresponding property of $(P_{t})_{t\geq 0}$. It remains to first let $t\to\infty$ and then $n\to\infty$ in (12). ∎
We will also need the following technical lemma.
Lemma 2.5.
Let $E$ be a Banach space and $F\in\mathbb{D}^{1,p}(E)$ for some $p\in[1,\infty)$. Then,
$\left\lVert D_{E}L_{E}^{-1}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}\leq\left\lVert D_{E}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}.$
Proof.
By density, it suffices to prove the claim for $F\in L^{p}(\Omega)\otimes E$. Applying  [Maa10, Lemma 6.2.i], we obtain $P_{E,t}F\in\mathbb{D}^{1,p}(E)$ and $D_{E}P_{E,t}F=e^{-t}P_{\gamma(\mathfrak{H},E),t}D_{E}F$, so that in particular
(13) $\left\lVert D_{E}P_{E,t}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}\leq e^{-t}\left\lVert D_{E}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}.$
Therefore,
$D_{E}L_{E}^{-1}F$ $=D_{E}\int_{0}^{\infty}P_{E,t}(F-\mathbb{E}(F))\,dt$
$=\int_{0}^{\infty}D_{E}P_{E,t}(F-\mathbb{E}(F))\,dt$
$=\int_{0}^{\infty}e^{-t}P_{\gamma(\mathfrak{H},E),t}D_{E}(F-\mathbb{E}(F))\,dt,$
with the interchange of integration and taking the Malliavin derivative being justified by (13), which furthermore implies
$\left\lVert D_{E}L_{E}^{-1}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}$ $\leq\int_{0}^{\infty}\left\lVert D_{E}P_{E,t}(F-\mathbb{E}(F))\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}$
$\leq\int_{0}^{\infty}e^{-t}dt\left\lVert D_{E}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}$
$=\left\lVert D_{E}F\right\rVert_{L^{p}(\Omega;\gamma(\mathfrak{H},E))}.$
Note that by construction, we have a natural correspondence between vector-valued Malliavin operators constructed so far and their classical scalar counterparts which they generalize. If $\Phi_{E}\in\left\{D_{E},P_{t,E},L_{E},L^{-1}_{E}\right\}$ is one of these operators with domain $\operatorname{dom}_{p}(\Phi_{E})\subseteq L^{p}(\Omega;E)$, it holds that
$\operatorname{dom}_{p}(\Phi_{\mathbb{R}})\otimes E\subseteq\operatorname{dom}_{p}(\Phi_{E}),$
where the inclusion is dense, and on $\operatorname{dom}_{p}(\Phi_{\mathbb{R}})\otimes E$ one has
$\Phi_{E}=\Phi_{\mathbb{R}}\otimes\operatorname{Id}_{E}.$
This implies that $\Phi_{E}$ commutes with linear forms $\ell\in E^{\ast}$, in the sense that if $F\in\operatorname{dom}_{p}(\Phi_{E})$, then $\ell(F)\in\operatorname{dom}_{p}(\Phi_{\mathbb{R}})$ and
$\Phi_{\mathbb{R}}\,\ell(F)=\ell(\Phi_{E}F).$
For this reason, we will most of the time refrain from adding the subscript $E$ to the respective operator from now on.
As another convention, whenever we use a Malliavin operator or consider a random element in its domain, we tacitly assume the existence of an underlying isonormal Gaussian process $W$ on a real, separable Hilbert space $\mathfrak{H}$, with respect to which the operator is defined.
2.7. Wiener-Itô chaos and multiple integrals in Banach spaces
Let $W\colon\mathfrak{H}\to L^{2}(\Omega,\mathcal{F},P)$ be an isonormal Gaussian process and $\mathcal{H}_{p}(\mathbb{R})$ the $p$-th Wiener chaos ($p\in\mathbb{N}_{0}$, with $\mathcal{H}_{0}(\mathbb{R})=\mathbb{R}$). We assume throughout this section that $\mathcal{F}$ coincides with the $\sigma$-field generated by $\left\{W(h)\colon h\in\mathfrak{H}\right\}$, so that the Wiener chaos decomposition
$L^{2}(\Omega,\mathcal{F},P)=\bigoplus_{p=0}^{\infty}\mathcal{H}_{p}(\mathbb{R})$
holds, where the sum is orthogonal and $\mathcal{F}$ is the $\sigma$-field generated by $W$.
For a Banach space $E$, the $E$-valued $p$-th Wiener chaos $\mathcal{H}_{p}(E)$ is defined as the $L^{2}(\Omega;E)$-completion of $\mathcal{H}_{p}(\mathbb{R})\otimes E$, the $E$-valued random elements which are linear combinations of $E$-vectors using coefficients in $\mathcal{H}_{p}(\mathbb{R})$. It follows from the Kahane-Khintchine inequality that, as is the case for $\mathcal{H}_{p}(\mathbb{R})$, all $L^{s}(\Omega;E)$-norms ($s\in[1,\infty)$) are equivalent on $\mathcal{H}_{p}(E)$.
As $\mathcal{F}_{sym}^{p}(\mathfrak{H},E)\simeq\mathfrak{H}^{\odot p}\otimes E$, we can define the tensor operator
$I_{p}^{\mathbb{R}}\otimes\operatorname{Id}_{E}\colon\mathcal{F}_{sym}^{p}(\mathfrak{H},E)\to\mathcal{H}_{p}(\mathbb{R})\otimes E,$
where $I_{p}^{\mathbb{R}}\colon\mathfrak{H}^{\odot p}\to\mathcal{H}_{p}(\mathbb{R})$ is the Wiener-Itô integral acting on real valued kernels.
It is one of the main findings of [Maa10] (see Thm. 3.2 therein), that for any $p\in\mathbb{N}_{0}$ and $s\in[1,\infty]$, the above tensor operator $I_{p}\otimes\operatorname{Id}_{E}$ extends to a bounded linear operator
$I_{p}^{E}\colon\gamma^{p}_{sym}(\mathfrak{H},E)\to L^{s}(\Omega;E),$
called the $E$-valued multiple Wiener-Itô integral of order $p$, whose range is $\mathcal{H}_{p}(E)$. The isometry property of $I_{p}^{\mathbb{R}}$ is only retained in the form of an equivalence of norms: For any $s\in[1,\infty)$ and any $p\in\mathbb{N}_{0}$ there exist positive constants $\mathfrak{c}_{p,s}$ and $\mathfrak{C}_{p,s}$ such that for all $T\in\gamma^{p}_{sym}(\mathfrak{H},E)$ one has
(14) $\mathfrak{c}_{p,s}\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)}\leq\left\lVert I_{p}^{E}(T)\right\rVert_{L^{s}(\Omega;E)}\leq\mathfrak{C}_{p,s}\left\lVert T\right\rVert_{\gamma^{p}(\mathfrak{H},E)},$
Inspecting the proof of [Maa10, Thm. 3.2] and in view of  [AG93, Thm. 2.2], one can take
(15) $\mathfrak{c}_{p,s}=2^{\frac{p-1}{s}}\sqrt{\frac{p!}{p^{p}}}\kappa_{s,2}^{-1}\qquad\text{and}\qquad\mathfrak{C}_{p,s}=4^{\frac{3p+1}{s}}\sqrt{\frac{p^{p}}{p!}}\kappa_{s,2},$
where $\kappa_{s,t}$ denotes the Kahane-Khintchine constant for the $L^{s}-L^{t}$-norm comparison of Gaussian (or Rademacher) sums (see for example [HvNVW17, Thm. 6.2.6] for a precise definition). In the special case where $E$ is a Hilbert space and $s=2$, (14) of course reduces to the classical Wiener-Itô isometry, i.e. the inequalities (14) become equalities, which hold for $\mathfrak{c}_{p,2}=\mathfrak{C}_{p,2}=\sqrt{p!}$.
Observe that, as for the other Malliavin operators (see the last paragraph in the previous section), the multiple integral commutes with linear forms in the following sense: if $f\in\gamma^{p}(\mathfrak{H},E)$ and $\ell\in E^{\ast}$, then $\ell\circ f\in\gamma^{p}(\mathfrak{H},\mathbb{R})\simeq\mathfrak{H}^{\widehat{\odot}_{\sigma}p}$ and
(16) $I_{p}^{\mathbb{R}}(\ell\circ f)=\ell\left(I_{p}^{E}(f)\right).$
Because of this compatibility result, we will most of the time refrain from adding the codomain of the kernel as a superscript to the operator $I_{p}$ in later sections. Taking the $L^{2}(\Omega)$-norm in (16) and applying the Wiener-Itô isometry for real-valued kernels gives
$\left\lVert\ell\left(I_{p}^{E}(f)\right)\right\rVert_{L^{2}(\Omega)}=\sqrt{p!}\left\lVert\ell\circ f\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}}.$
One can verify directly that kernels $f\in\mathfrak{H}^{\widehat{\odot}_{\sigma}p}\widehat{\otimes}_{\pi}E$ of the form
$f=\sum_{j=1}^{\infty}f_{j}\otimes x_{j}$
with $f_{j}\in\mathfrak{H}^{\widehat{\odot}_{\sigma}p}$ and $x_{j}\in E$ for $j\in\mathbb{N}$ and such that
$\sum_{j=1}^{\infty}\left\lVert f_{j}\right\rVert_{\mathfrak{H}^{\widehat{\odot}_{\sigma}p}}\left\lVert x_{j}\right\rVert_{E}<\infty$
are in $\gamma^{p}_{sym}(\mathfrak{H},E)$ (it suffices to approximate in $\mathcal{H}^{\odot p}\otimes E$ and crudely bound the $\gamma^{p}(\mathfrak{H},E)$-norm by shifting the $L^{2}$-norm to the summands).
By the Wiener-Itô isometry, the corresponding integrals are nuclear series of the form
$I_{p}^{E}(f)=\sum_{j=1}^{\infty}I_{p}^{\mathbb{R}}(f_{j})\otimes x_{j}$
with
$\sum_{j=1}^{\infty}\left\lVert f_{j}\right\rVert_{\mathfrak{H}^{\widehat{\odot}_{\sigma}p}}\left\lVert x_{j}\right\rVert_{E}<\infty.$
Such integrals, in particular finite-rank versions where all sums are finite, will play a role in the applications treated in later sections.
When generalizing the multiple integral for $\mathfrak{H}=L^{2}(M)=L^{2}(M,\mathcal{M},\mu)$, with $\mu$ a $\sigma$-finite measure without atoms on a measurable space $(M,\mathcal{M})$, some care has to be applied as one can no longer use the isometric identification
$L^{2}(M,\mathcal{M},\mu)^{\widehat{\otimes}_{\sigma}p}\simeq L^{2}(M^{p},\mathcal{M}^{\otimes p},\mu^{\otimes p}).$
Indeed, as mentioned earlier, $\gamma^{p}(\mathfrak{H},E)$ is in general not isomorphic to $\gamma(\mathfrak{H}^{\widehat{\otimes}_{\sigma}p},E)$. However, as noted in [Maa10, Section 2.2], the space $\gamma^{p}(\mathfrak{H},E)$ contains a dense subset of representable operators $f$ for which there exists a Bochner measurable function $\varphi_{f}\colon M^{p}\to E$ which is weakly $L^{2}$, i.e. $(\varphi,x^{\ast})_{E,E^{\ast}}\in L^{2}(\mu^{\otimes p})$ for all $x^{\ast}\in E^{\ast}$, such that
$(f(h_{1},h_{2},\dots,h_{p}),x^{\ast})_{E,E^{\ast}}=\int_{M^{p}}h_{1}(t_{1})\dots h_{p}(t_{p})\left(\phi_{f}(t_{1},\dots,t_{p}),x^{\ast}\right)_{E,E^{\ast}}d\mu^{\otimes p}\left(t_{1},\dots,t_{p}\right).$
For such representable operators, the multiple integral can be lifted to the vector-valued case (see [Maa10, Sec. 4]).
Let us point out that the tensor operators $J_{p}\otimes\operatorname{Id}_{E}$, where $J_{p}$ is the bounded projection onto the $p$-th Wiener chaos $\mathcal{H}_{p}(\mathbb{R})$, do in general not extend to bounded projections onto $\mathcal{H}_{p}(E)$. The needed property here is $K$-convexity of $E$, possessed for example by UMD spaces and in particular Hilbert spaces. See [Maa10, Rem. 3.4] for references and details.
The projection $J_{0}$ onto the constants $\mathcal{H}_{0}(E)$ however is always bounded, as it coincides with the expectation operator (see the construction of $L^{-1}$ in the previous section).
3. The carré du champ operator and integration by parts formulæ
In this section, we construct a tensor-valued carré du champ operator and then use it to prove integration by parts formulæ, which can be interpreted as chain rules for (cross)-covariance tensors and operators. Applying the smart path method, these chain rules yield a generic carré du champ-bound for differences of the form
$\left|\mathbb{E}\left(f(X)\right)-\mathbb{E}\left(f(Y)\right)\right|,$
with sufficiently regular Banach-valued random elements $X$ and $Y$ and test functions $f$, which will serve as a prototype to bound integral probability metrics in the later sections.
We begin by introducing a tensor contraction, also known as vector valued trace (see  [DF93, Ch. 29] for a more general definition and the associated calculus).
Let $\mathfrak{H}$ be a pre-Hilbert space and $E$, $F$ be normed spaces. The contraction $C_{\mathfrak{H}}\colon(\mathfrak{H}\otimes E)\times(\mathfrak{H}\otimes F)\to E\otimes F$ is defined for elementary tensors as
$C_{\mathfrak{H}}(h\otimes x,h^{\prime}\otimes y)=\left\langle h,h^{\prime}\right\rangle_{\mathfrak{H}}x\otimes y$
and then extended by bilinearity. If the pre-Hilbert space is of the form $\mathfrak{H}^{\otimes r}$, we write $C_{\mathfrak{H}}^{r}=C_{\mathfrak{H}^{\otimes r}}$ (so that $C_{\mathfrak{H}}^{1}=C_{\mathfrak{H}}$), and in addition define
$C_{\mathfrak{H}}^{0}\colon(\mathfrak{H}\otimes E)\times(\mathfrak{H}\otimes F)$ $\to\mathfrak{H}\otimes\mathfrak{H}\otimes E\otimes F$
$(h\otimes x,h^{\prime}\otimes y)$ $\mapsto h\otimes h^{\prime}\otimes x\otimes y.$
Canonically identifying the algebraic tensor product $\mathfrak{H}^{\otimes p}\otimes E$ with the finite-rank operators $\mathcal{F}^{p}(\mathfrak{H};E)$, the contraction is also defined as an operator
$C_{\mathfrak{H}}^{r}\colon\mathcal{F}^{p}(\mathfrak{H};E)\times\mathcal{F}^{p}(\mathfrak{H};F)\to\mathcal{F}^{p+q-2r}(\mathfrak{H};E\otimes F)$
for any $p,q\in\mathbb{N}$ and $0\leq r\leq p\land q$, and of course we could further identify the algebraic tensor product $E\otimes F$ with the finite rank operators $\mathcal{F}(F^{\ast},E)$. For a normed space $S$, we denote by $\mathcal{F}_{\gamma}(\mathfrak{H},S)$ the space $\mathcal{F}(\mathfrak{H},S)$ equipped with the gamma-radonifying norm $\gamma$ and then inductively define $\mathcal{F}^{p}_{\gamma}(\mathfrak{H},S)$. The following lemma shows that the contraction has a continuous extension to spaces of $\gamma$-radonifying operators (see Section 2.5 for the definition).
Lemma 3.1.
For $p,q\in\mathbb{N}$, a real separable Hilbert space $\mathfrak{H}$ and Banach spaces $E$, $F$, let $f\in\mathcal{F}^{p}(\mathfrak{H},E)$ and $g\in\mathcal{F}^{q}(\mathfrak{H},F)$. Then, for any $0\leq r\leq p\land q$, one has
(17) $\left\lVert C_{\mathfrak{H}}^{r}(f,g)\right\rVert_{\mathcal{F}^{p+q-2r}_{\gamma}(\mathfrak{H},E\otimes_{\pi}F)}\leq\left\lVert f\right\rVert_{\mathcal{F}^{p}_{\gamma}(\mathfrak{H},E)}\left\lVert g\right\rVert_{\mathcal{F}^{q}_{\gamma}(\mathfrak{H},F)}.$
Consequently, the contraction operator has a bounded tensor-valued extension
$\widehat{C}_{\mathfrak{H},\pi}^{r}\colon\gamma^{p}(\mathfrak{H},E)\times\gamma^{q}(\mathfrak{H},F)\to\gamma^{p+q-2r}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)$
and also a bounded operator-valued extension
$\widehat{C}_{\mathfrak{H}}^{r}\colon\gamma^{p}(\mathfrak{H},E)\times\gamma^{q}(\mathfrak{H},F)\to\gamma^{p+q-2r}(\mathfrak{H},\mathfrak{N}(F^{\ast},E))$
Remark 3.2.
  1. (i)
    Note that in the case $r=1$, we have $\widehat{C}_{\mathfrak{H}}(f,g)=fg^{\ast}$, where $g^{\ast}$ denotes the adjoint of $g$.
  2. (ii)
    Lemma 3.1 also yields bounded versions of $C_{\mathfrak{H}}^{r}$ mapping into $\gamma^{p+q-2r}(\mathfrak{H},E\widehat{\otimes}_{\alpha}F)$ for any tensor norm $\alpha$, as $\alpha\leq\pi$. A similar statement can be made for the associated operator ideals.
Proof of Lemma 3.1.
[…]
$\left\lVert T_{u}\right\rVert_{\nu}\leq\left\lVert u\right\rVert_{\pi}.$
Consequently, identifying $T\in\mathcal{F}^{p}_{\gamma}(\mathfrak{H},E\otimes_{\pi}F)$ with $\widetilde{T}\in\mathcal{F}^{p}_{\gamma}(\mathfrak{H},\mathcal{F}_{\nu}(F^{\ast},E))$, monotonicity of the expectation yields
$\left\lVert\widetilde{T}\right\rVert_{\mathcal{F}^{p}_{\gamma}(\mathfrak{H},\mathcal{F}_{\nu}(F^{\ast},E))}\leq\left\lVert T\right\rVert_{\mathcal{F}^{p}_{\gamma}(\mathfrak{H},E\otimes_{\pi}F)},$
where $\mathcal{F}_{\nu}(F^{\ast},E)$ denotes the finite rank operators equipped with the nuclear norm, which are dense in $\mathfrak{N}(F^{\ast},E)$. Together with (17), we get
$\left\lVert C_{\mathfrak{H}}^{r}(f,g)\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},\mathcal{F}^{(\nu)}(F^{\ast},E))}\leq\left\lVert f\right\rVert_{\mathcal{F}^{p}_{\gamma}(\mathfrak{H},E)}\left\lVert g\right\rVert_{\mathcal{F}^{q}_{\gamma}(\mathfrak{H},F)}$
for $f\in\mathcal{F}^{p}_{\gamma}(\mathfrak{H},E)$ and $g\in\mathcal{F}^{q}_{\gamma}(\mathfrak{H},F)$ and can then pass to the completion.
To prove (17), we will reuse the multi index notation introduced in Section 2.5.
Fix an orthonormal basis $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ of $\mathfrak{H}$ and let $f\in\mathcal{F}^{p}(\mathfrak{H},E)$, $g\in\mathcal{F}^{p}(\mathfrak{H},F)$ with representations analogous to (9) in Section 2.5 of the form
$f=\sum_{\mathbf{i}\in[n]^{p}}\mathfrak{h}_{\mathbf{i}}\otimes x_{\mathbf{i}}\qquad\text{and}\qquad g=\sum_{\mathbf{j}\in[n]^{q}}\mathfrak{h}_{\mathbf{j}}\otimes y_{\mathbf{j}}$
with $x_{\mathbf{i}}\in E$ and $y_{\mathbf{j}}\in F$ for all $\mathbf{i}\in[n]^{p}$ and $\mathbf{j}\in[n]^{q}$, where $[n]$ denotes the set of the first $n$ integers and $n]^{p}$ the corresponding $p$-tuples thereof. The notation $\mathfrak{h}_{\mathbf{j}}\otimes y_{\mathbf{j}}$ for a rank-one operator retains the same meaning as in (9), so does $W_{\mathbf{i}}$ for the product of $m$ independent copies of the isonormal Gaussian process $W$ evaluated in $\mathfrak{h}_{j}$. Let $^{\prime}, W′′W^{\prime\prime}$ be independent copies of $W$ as well and denote by $\mathbb{E}^{\prime}$ and $\mathbb{E}^{\prime\prime}$ expectation with respect to $W^{\prime}$ and $W^{\prime\prime}$, respectively, including all its respective independent copies if we use the multi-index notation $W^{\prime}_{\mathbf{i}}$ or $W^{\prime\prime}_{\mathbf{i}}$), whereas the expectation $\mathbb{E}$ is always taken with respect to all random elements inside its argument. Finally, for two multi-indices $\mathbf{k}\in\mathbb{N}^{m}$ and $\mathbf{l}\in\mathbb{N}^{n}$, we denote their concatenation $(k_{1},\dots,k_{m},l_{1},\dots,l_{n})\in\mathbb{N}^{m+n}$ by $(\mathbf{k},\mathbf{l})$.
Fix $1\leq r\leq p\land q$. By definition of the contraction operator and the $\gamma$-radonifying norm, we obtain
$\left\lVert C_{\mathfrak{H}}^{r}(f,g)\right\rVert^{2}_{\gamma^{p+q-2r}(\mathfrak{H},E\otimes_{\pi}F)}$
$\qquad=\mathbb{E}\left(\left\lVert\sum_{\begin{subarray}{c}\mathbf{u}\in[n]^{r}\\ \mathbf{s}\in[n]^{p-r},\mathbf{t}\in[n]^{q-r}\end{subarray}}\left(W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right)\otimes\left(W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right)\right\rVert_{\pi}^{2}\right)$
$\qquad=\mathbb{E}\left(\left\lVert\mathbb{E}^{\prime\prime}\left(\sum_{\begin{subarray}{c}\mathbf{u},\mathbf{v}\in[n]^{r}\\ \mathbf{s}\in[n]^{p-r},\mathbf{t}\in[n]^{q-r}\end{subarray}}\left(W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right)\otimes\left(W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right)\right)\right\rVert_{\pi}^{2}\right)$
$\qquad=\mathbb{E}\left(\left\lVert\mathbb{E}^{\prime\prime}\left(\left(\sum_{\mathbf{u}\in[n]^{r},\mathbf{s}\in[n]^{p-r}}W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right)\otimes\left(\sum_{\mathbf{v}\in[n]^{r},\mathbf{t}\in[n]^{q-r}}W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right)\right)\right\rVert_{\pi}^{2}\right),$
where we have used Gaussian orthogonality. By monotonicity of the Bochner integral and the cross norm property of the projective norm ($\left\lVert x\otimes y\right\rVert_{\pi}=\left\lVert x\right\rVert_{E}\cdot\left\lVert y\right\rVert_{F}$), we can continue to write
$\qquad\leq\mathbb{E}\left(\mathbb{E}^{\prime\prime}\left(\left\lVert\left(\sum_{\mathbf{u}\in[n]^{r},\mathbf{s}\in[n]^{p-r}}W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right)\otimes\left(\sum_{\mathbf{v}\in[n]^{r},\mathbf{t}\in[n]^{q-r}}W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right)\right\rVert_{\pi}\right)^{2}\right)$
$\qquad=\mathbb{E}\left(\mathbb{E}^{\prime\prime}\left(\left\lVert\sum_{\mathbf{u}\in[n]^{r},\mathbf{s}\in[n]^{p-r}}W^{\prime\prime}_{\mathbf{u}}W_{\mathbf{s}}\,x_{(\mathbf{u},\mathbf{s})}\right\rVert_{E}\,\left\lVert\sum_{\mathbf{v}\in[n]^{r},\mathbf{t}\in[n]^{q-r}}W^{\prime\prime}_{\mathbf{v}}W^{\prime}_{\mathbf{t}}\,y_{(\mathbf{u},\mathbf{t})}\right\rVert_{F}\right)^{2}\right),$
and it remains to apply Cauchy-Schwarz to bound the above by $\left\lVert f\right\rVert_{\gamma^{p}(\mathfrak{H},E)}^{2}\,\left\lVert g\right\rVert_{\gamma^{q}(\mathfrak{H},F)}^{2}$.
The proof for the case $r=0$ follows in a similar but simpler way. ∎
We can now introduce the carré du champ operator.
Definition 3.3.
Let $E$ and $F$ be Banach spaces. The tensor valued carré du champ operator $\Gamma_{\pi}\colon\mathbb{D}^{1,2}(E)\times\mathbb{D}^{1,2}(F)\to L^{1}(\Omega;E\widehat{\otimes}_{\pi}F)$ is defined by
$\Gamma_{\pi}(X,Y)=\widehat{C}_{\mathfrak{H},\pi}(DX,DY),$
and the operator valued carré du champ operator $\Gamma\colon\mathbb{D}^{1,2}(E)\times\mathbb{D}^{1,2}(F)\to L^{1}(\Omega;\mathfrak{N}(F^{\ast},E)$ as
$\Gamma(X,Y)=\widehat{C}_{\mathfrak{H}}(DX,DY)=DX(DY)^{\ast},$
where $\widehat{C}_{\mathfrak{H},\pi}$ and $\widehat{C}_{\mathfrak{H}}$ are the bounded contraction operators defined in Lemma 3.1.
Lemma 3.1 immediately yields the boundedness of both carré du champ operators, as
$\mathbb{E}\left(\left\lVert\Gamma(X,Y)\right\rVert_{\nu}\right)\leq\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,Y)\right\rVert_{\pi}\right)=\mathbb{E}\left(\left\lVert\widehat{C}_{\mathfrak{H},\pi}(DX,DY)\right\rVert_{\pi}\right)\\ \leq\mathbb{E}\left(\left\lVert DX\right\rVert_{\gamma(\mathfrak{H},E)}\left\lVert DY\right\rVert_{\gamma(\mathfrak{H},F)}\right)\leq\left\lVert DX\right\rVert_{1,2}\left\lVert DY\right\rVert_{1,2},$
where $\left\lVert\cdot\right\rVert_{1,2}$ denotes the norm of $\mathbb{D}^{1,2}(E)$. Arguing analogously to Remark 3.2, the tensor valued carré du champ operator is also bounded as a random bilinear mapping taking values in $L^{1}(\Omega;E\widehat{\otimes}_{\alpha}F)$, where $\alpha$ is any tensor norm.
The chain rule for the Malliavin derivative (Theorem 2.3) immediately yields the following chain rule for the carré du champ.
Lemma 3.4.
Let $E_{1}$, $E_{2}$, $F_{1}$, $F_{2}$ be Banach spaces, $X\in\mathbb{D}^{1,2}(E_{1})$, $Y\in\mathbb{D}^{1,2}(F_{1})$, $f\in\mathcal{C}^{1}_{b}(E_{1},E_{2})$ and $g\in\mathcal{C}^{1}_{b}(F_{1},F_{2})$.
  1. (i)
    If $E_{2}$ has the UMD property, then
    $\Gamma\left(f(X),Y\right)=f^{\prime}(X)\Gamma(X,Y)$
    and
    $\Gamma_{\pi}\left(f(X),Y\right)=\left(f^{\prime}(X)\otimes\operatorname{Id}_{F_{1}}\right)\Gamma_{\pi}\left(X,Y\right)$
  2. (ii)
    If $F_{2}$ has the UMD property, then
    $\Gamma\left(X,g(Y)\right)=\Gamma\left(X,Y\right)g(Y)^{\ast}$
    and
    $\Gamma_{\pi}\left(X,g(Y)\right)=\left(\operatorname{Id}_{E_{1}}\otimes g^{\prime}(Y)\right)\Gamma_{\pi}\left(X,Y\right).$
The next lemma will be crucial for obtaining the integration by parts formula.
Lemma 3.5.
Let $E$ and $F$ be Banach spaces, $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$. Then,
(18) $Q_{X,Y}=\mathbb{E}\left(\Gamma_{\pi}\left(X,-L^{-1}Y\right)\right)=\mathbb{E}\left(\Gamma_{\pi}\left(-L^{-1}X,Y\right)\right)$
and
(19) $R_{X,Y}=\mathbb{E}\left(\Gamma\left(X,-L^{-1}Y\right)\right)=\mathbb{E}\left(\Gamma\left(-L^{-1}X,Y\right)\right).$
Proof.
By Lemma 2.5 and Lemma 3.1, the carré du champ remains continuous as a bilinear mapping from $\mathbb{D}^{1,2}(E)\times\mathbb{D}^{1,2}(F)$ to $L^{1}(\Omega;E\widehat{\otimes}_{\pi}F)$ (or $L^{1}(\Omega;\mathfrak{N}(F^{\ast},E))$) when we apply $L^{-1}$ to one of its arguments. Approximating in $\mathbb{D}^{1,2}(E)$ and $\mathbb{D}^{1,2}(F)$, it suffices to show the identities (18) for $X$ and $Y$ in the dense subsets $\mathcal{S}(\mathbb{R})\otimes E$ and $\mathcal{S}(\mathbb{R})\otimes F$, respectively. For such $X$ and $Y$, write
$X=\sum_{j=1}^{n}X_{j}\otimes x_{j}\qquad\text{and}\qquad Y=\sum_{k=1}^{n}Y_{k}\otimes y_{k}$
with $X_{j},Y_{k}\in\mathbb{D}^{1,2}(\mathbb{R})$, $x_{j}\in E$ and $y_{k}\in F$ for $1\leq k,l\leq n$ and some $n\in\mathbb{N}$. Then,
(20) $Q_{X,Y}=\mathbb{E}\left(\left(X-\mathbb{E}(X)\right)\otimes Y\right)=\sum_{j,k=1}^{n}\mathbb{E}\left(\left(X_{j}-\mathbb{E}(X_{j})\right)Y_{k}\right)x_{j}\otimes y_{k}=\sum_{j,k=1}^{n}R_{X_{j},Y_{k}}\,x_{j}\otimes y_{k}.$
Let $\mathfrak{H}$ be the real, separable Hilbert space on which the underlying isonormal Gaussian process is defined and pick an orthonormal basis $\left\{\mathfrak{h}_{n}\colon n\in\mathbb{N}\right\}$. As both $X_{j}-\mathbb{E}(X_{j})$ and $Y_{k}$ are real valued, the well-known scalar-valued integration by parts formula yields
$R_{X_{j},Y_{k}}=\mathbb{E}\left(\left(X_{j}-\mathbb{E}(X_{j})\right)\left(Y_{k}-\mathbb{E}(Y_{k})\right)\right)=-\mathbb{E}\left(X_{j}\left(LL^{-1}Y_{k}\right)\right)=\mathbb{E}\left(\Gamma(X_{j},-L^{-1}Y_{k})\right).$
Plugged into (20), this yields
$Q_{X,Y}$ $=\sum_{j,k=1}^{n}\mathbb{E}\left(\Gamma(X_{j},-L^{-1}Y_{k})\right)x_{j}\otimes y_{k}$
$=\sum_{j,k=1}^{n}\mathbb{E}\left(\sum_{l=1}^{\infty}\left(DX_{j},\mathfrak{h}_{l}\right)_{\mathfrak{H}}\left(-DL^{-1}Y_{k},\mathfrak{h}_{l}\right)_{\mathfrak{H}}\right)\,x_{j}\otimes y_{k}$
$=\sum_{j,k=1}^{n}\mathbb{E}\left(\sum_{l=1}^{\infty}\left(D(X_{j}\otimes x_{j}),\mathfrak{h}_{l}\right)_{\mathfrak{H}}\left(-DL^{-1}(Y_{k}\otimes y_{k}),\mathfrak{h}_{l}\right)_{\mathfrak{H}}\right)$
$=\sum_{j,k=1}^{n}\mathbb{E}\left(\Gamma\left(X_{j}\otimes x_{j},-L^{-1}\left(Y_{k}\otimes y_{k}\right)\right)\right)$
$=\mathbb{E}\left(\Gamma\left(X,-L^{-1}Y\right)\right),$
and, as $R_{X_{j},Y_{k}}=R_{Y_{k},X_{j}}$, also
$Q_{X,Y}=\mathbb{E}\left(\Gamma\left(-L^{-1}X,Y\right)\right).$
The identity (19) for the operator-valued carré du champ follows by identifying the algebraic tensors with the respective finite-rank operators. ∎
We are now ready to prove that the integration by parts formula continues to hold in the Banach-valued setting. Recall the discussion on topological tensor products in Section 2.2.
Proposition 3.6 (Malliavin integration by parts for Banach-valued random elements).
For Banach spaces $E$ and $F$, let $X\in\mathbb{D}^{1,2}(E)$, $Y\in\mathbb{D}^{1,2}(F)$ and $f\in\mathcal{C}^{1}_{b}(E,F^{\ast})$. Then,
$\operatorname{tr}_{\pi}\left(Q_{f(X),Y}\right)=\mathbb{E}\left(\left(f^{\prime}(X),\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right),$
where $f^{\prime}(x)$ is identified with an element of $\mathcal{L}(E,F;\mathbb{R})$ and $\left(\cdot,\cdot\right)_{\pi}$ is the pairing between the projective tensor product and its dual. If $F$ has the approximation property, we also have that
$\operatorname{tr}\left(R_{f(X),Y}\right)=\operatorname{tr}\left(\mathbb{E}\left(f^{\prime}(X)\Gamma(X,-L^{-1}Y)\right)\right).$
Proof.
We first treat the case where $X$ and $Y$ are of the form
$X=\sum_{j=1}^{n}X_{j}\otimes x_{j}\qquad\text{and}\qquad Y=\sum_{k=1}^{n}Y_{k}\otimes y_{k}$
with $X_{j},Y_{k}\in\mathcal{S}(\mathbb{R})$, $x_{j}\in E$ and $y_{k}\in F$ for $1\leq k,l\leq n$ and some $n\in\mathbb{N}$. Let $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ be an orthonormal basis of $\mathfrak{H}$. Then,
$\mathbb{E}\left(\left(f(X),Y-\mathbb{E}(Y)\right)_{F^{\ast},F}\right)$ $=\sum_{k=1}^{n}\mathbb{E}\left(\left(Y_{k}-\mathbb{E}(Y_{k})\right)\left(f(X),y_{k}\right)_{F^{\ast},F}\right)$
(21) $=\sum_{k=1}^{n}\mathbb{E}\left(\left(Y_{k}-\mathbb{E}(Y_{k})\right)\varphi_{k}(X)\right),$
where $\varphi_{k}\in\mathcal{C}_{b}^{1}(E,\mathbb{R})$ is defined as $\varphi_{k}(x)=\left(f(x),y_{k}\right)_{F^{\ast},F}$ for $k=1,\dots,n$. The Fréchet derivative of $\varphi_{k}$ is given by
$\varphi^{\prime}_{k}(x)(u)=\left(f^{\prime}(x)u,y_{k}\right)_{F^{\ast},F}$
for $x,u\in E$, so that by the chain rule for the Malliavin derivative (Theorem 2.3),
$D\varphi_{k}(X)=\varphi_{k}^{\prime}(X)DX=\left(f^{\prime}(X)DX,y_{k}\right)_{F^{\ast},F}.$
Furthermore, the well-known finite-dimensional integration by parts formula gives
$\mathbb{E}\left(\left(Y_{k}-\mathbb{E}(Y_{k})\right)\varphi_{k}(X)\right)$ $=\mathbb{E}\left(\left(D\varphi_{k}(X),-DL^{-1}Y_{k}\right)_{\mathfrak{H}}\right)$
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},y_{k}\right)_{F^{\ast},F}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)\right)$
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)y_{k}\right)_{F^{\ast},F}\right),$
where only finitely many summands are different from zero as the finitely many $X_{j}$ and $Y_{k}$ individually and hence collectively only depend on finitely many Gaussians $W(\mathfrak{h}_{l})$ (with $W$ denoting the underlying isonormal Gaussian process). Plugged into (21), writing $b_{f^{\prime}(X)}$ for the bounded bilinear form on $F$ determined by $f^{\prime}(X)$ via $b_{f^{\prime}(X)}(u,v)=f^{\prime}(X)(u)(v)$, we obtain
$\operatorname{tr}_{\pi}\left(Q_{f(X),Y}\right)$ $=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},\sum_{k=1}^{n}\left(-DL^{-1}Y_{k}\mathfrak{h}_{l}\right)y_{k}\right)_{F^{\ast},F}\right)$
$=\mathbb{E}\left(\sum_{l}\left(f^{\prime}(X)DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)_{F^{\ast},F}\right)$
$=\mathbb{E}\left(\sum_{l}b_{f^{\prime}(X)}\left(DX\,\mathfrak{h}_{l},-DL^{-1}Y\mathfrak{h}_{l}\right)\right)$
$=\mathbb{E}\left(\left(b_{f^{\prime}(X)},\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right).$
It is clear that we also have
$\operatorname{tr}\left(R_{f(X),Y}\right)=\operatorname{tr}\left(\mathbb{E}\left(f^{\prime}(X)\Gamma(X,-L^{-1}Y)\right)\right)=\mathbb{E}\left(\operatorname{tr}\left(f^{\prime}(X)\Gamma(X,-L^{-1}Y)\right)\right),$
where the second equality follows from the fact that the Bochner integral commutes with bounded linear mappings.
For the general case, we approximate $X\in\mathbb{D}^{1,2}(E)$ and $Y\in\mathbb{D}^{1,2}(F)$ by sequences $(X_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}\left(\mathcal{S}(\mathbb{R}),E\right)$ and $(Y_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}\left(\mathcal{S}(\mathbb{R}),F\right)$, such that $\left\lVert X-X_{n}\right\rVert_{1,2}\to 0$ and $\left\lVert Y-Y_{n}\right\rVert_{1,2}\to 0$. Possibly passing to subsequences, we assume that $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ converge almost surely. Then,
$\operatorname{tr}_{\pi}\left(Q_{f(X_{n}),Y_{n}}\right)=\mathbb{E}\left(\left(b_{f^{\prime}(X_{n})},\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right).$
First, note that
$\operatorname{tr}_{\pi}$ $\left(Q_{f(X),Y}\right)-\operatorname{tr}_{\pi}\left(Q_{f(X_{n}),Y_{n}}\right)$
$=\mathbb{E}\left(\left|\left(Y-\mathbb{E}(Y),f(X)\right)_{F,F^{\ast}}-\left(Y_{n}-\mathbb{E}(Y_{n}),f(X_{n})\right)_{F,F^{\ast}}\right|\right)$
$\leq\mathbb{E}\left|\left(Y-\mathbb{E}(Y),f(X)-f(X_{n})\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(Y-Y_{n},f(X_{n})\right)_{F,F^{\ast}}\right|+\mathbb{E}\left|\left(\mathbb{E}(Y-Y_{n}),f(X_{n})\right)_{F,F^{\ast}}\right|$
$\leq 2\left\lVert f\right\rVert_{\infty}\mathbb{E}\left(\left\lVert Y-Y_{n}\right\rVert_{F}\right)+2\mathbb{E}\left(\left\lVert Y\right\rVert_{F}\right)\mathbb{E}\left(\left\lVert f(X)-f(X_{n})\right\rVert_{F^{\ast}}\right)$
which tends to zero as $n\to\infty$ by dominated convergence and the assumption that $Y_{n}\to Y$ almost surely. Furthermore
$\mathbb{E}$ $\left|\left(f^{\prime}(X),\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}-\left(f^{\prime}(X_{n}),\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right|$
$\qquad\leq\mathbb{E}\left|\left(f^{\prime}(X)-f^{\prime}(X_{n}),\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right|$
$\qquad\qquad\qquad+\mathbb{E}\left|\left(f^{\prime}(X_{n}),\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right)_{\pi}\right|$
$\qquad\leq\mathbb{E}\left(\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}\,\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)\right\rVert_{\pi}\right)$
$\qquad\qquad\qquad+\left\lVert f^{\prime}\right\rVert_{\infty}\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})\right\rVert_{\pi}\right)$
and both terms on the right hand side vanish in the limit due to dominated convergence, as $\left\lVert\Gamma_{\pi}(X_{n},-L^{-1}Y_{n})-\Gamma_{\pi}(X,-L^{-1}Y)\right\rVert_{\pi}$ converges to zero by the boundedness of $\Gamma_{\pi}$ and $\left\lVert f^{\prime}(X)-f^{\prime}(X_{n})\right\rVert_{\mathcal{L}(E,F;\mathbb{R})}$ converges to zero by continuity of $f$ and the almost sure convergence of $(X_{n})_{n\in\mathbb{N}}$.
If $F$ has the metric approximation property, the trace is defined as a bounded linear functional on the space of nuclear operators $\mathfrak{N}(F^{\ast},F)$, in which $f^{\prime}(X)\Gamma(X,-L^{-1}Y)$ takes its values, so that we can can complete the finite-rank operators with respect to the nuclear norm. In this case, the above calculation hence also yields the identity for the operator valued carré du champ. ∎
We now show that jointly Radon Gaussians can always be represented as Radonifications of a common underlying isonormal Gaussian process. Their cross-covariance operators then coincide almost surely with the corresponding carré du champ operators. This result is in complete analogy with the scalar-valued case.
Proposition 3.7.
Let $Z_{1}$ and $Z_{2}$ be jointly Radon Gaussians, taking values in Banach spaces $E_{1}$ and $E_{2}$, respectively. Then there exists a real, separable Hilbert space $\mathfrak{H}$, an isonormal Gaussian process $W\colon\mathfrak{H}\to L^{2}(\Omega)$ and $\gamma$-radonifying operators $T_{Z_{i}}\colon\mathfrak{H}\to E_{i}$, such that the law of $Z_{i}$ coincides with the Radonification of $W$ by $T_{Z_{i}}$ and
$R_{Z_{i},Z_{j}}=\Gamma\left(Z_{i},-L^{-1}Z_{j}\right)\qquad\text{and}\qquad Q_{Z_{i},Z_{j}}=\Gamma_{\pi}\left(Z_{i},-L^{-1}Z_{j}\right)$
for all $1\leq i,j\leq 2$, where all identities hold almost surely. In particular, the carré du champ of Radon Gaussians is almost surely deterministic.
Proof.
As for constant random elements $X$, one has $DX=L^{-1}X=0$ almost surely, it follows that
$\mathbb{E}\left(\Gamma\left(Z_{i},-DL^{-1}Z_{j}\right)\right)=\mathbb{E}\left(\Gamma(Z_{i}-\mathbb{E}(Z_{i}),-L^{-1}Z_{j})\right)=\mathbb{E}\left(\Gamma\left(Z_{i},-L^{-1}\left(Z_{j}-\mathbb{E}(Z_{j})\right)\right)\right),$
and the same identities are true for $\Gamma_{\pi}$. Therefore, it suffices to show the claim for centered Gaussians $Z_{1}$ and $Z_{2}$.
Let $G=E_{1}\oplus E_{2}$. It is straightforward to verify that the covariance operator $R_{Z}\colon G^{\ast}\to G$ of $Z=Z_{1}\oplus Z_{2}$ is given by the matrix $(R_{Z_{i},Z_{j}})_{i,j=1,2}$, by which we mean that
$R_{Z}(e_{1}^{\ast}\oplus e_{2}^{\ast})=\left(R_{Z_{1},Z_{1}}e_{1}^{\ast}+R_{Z_{1},Z_{2}}e_{2}^{\ast})\oplus(R_{Z_{2},Z_{1}}e_{1}^{\ast}+R_{Z_{2},Z_{2}}e_{2}^{\ast}\right)$
for all $e_{1}^{\ast}\oplus e_{2}^{\ast}\in G^{\ast}$.
Let $\mathfrak{H}\subseteq G$ be the Cameron-Martin space of the law of $Z$ and note that $\mathfrak{H}$ is separable as $Z$ is Radon. Also recall that $\mathfrak{H}$ is a vector subspace of $G$.
Choose an orthonormal basis $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ of $\mathfrak{H}$ for which
$\sum_{k=1}^{\infty}\left\lVert\mathfrak{h}_{k}\right\rVert_{E}^{2}<\infty.$
As mentioned in Section 2.4, such a basis always exists. Then, the $G$-valued Gaussian series
(22) $\sum_{k}^{\infty}\xi_{k}\mathfrak{h}_{k},$
where the $\xi_{k}$ are independent real-valued standard Gaussians on some probability space $(\Omega,\mathcal{F},P)$, converges almost surely and in $L^{s}(\Omega;G)$ for any $s\in[1,\infty)$ to a Gaussian random element with the same law as $Z$ (see for example [Bog98, Theorem 3.5.1]). Without loss of generality, we will take this series as the definition of $Z$ from now on. The series also defines an isonormal Gaussian process $W\colon\mathfrak{H}\to L^{2}(\Omega)$ by
$W=\sum_{k=1}^{\infty}\xi_{k}\left(\mathfrak{h}_{k},\cdot\right)_{\mathfrak{H}},$
for which the canonical embedding $i_{Z}\colon\mathfrak{H}\to G$ serves as a $\gamma$-radonifying operator, i.e. we have that $(f,Z)_{G^{\ast},G}=Wi_{Z}^{\ast}f$ for all $f\in G^{\ast}$, where we have identified $\mathfrak{H}$ with its dual. Note that $\mathfrak{H}$ here is at the same time the domain of the isonormal Gaussian process and the Cameron-Martin space of the Gaussian $Z$, while in general these two spaces do typically not coincide (but are isometrically isomorphic).
While square summability of the $G$-norms of the $\mathfrak{h}_{k}$ is not needed for the Gaussian series above to converge, it yields the nuclear representation
$R_{Z}=\mathbb{E}\left((\cdot,Z)_{G^{\ast},G}Z\right)=\sum_{k=1}^{\infty}(\cdot,i_{Z}\mathfrak{h}_{k})_{G^{\ast},G}i_{Z}\mathfrak{h}_{k}$
for its covariance operator.
For $i=1,2$, let $P_{i}\colon G\to E_{i}$ be the canonical bounded projection onto the closed subspace $E_{i}$. Then $T_{i}=P_{i}i_{Z}\colon\mathfrak{H}\to E_{i}$ is bounded (as the composition of two bounded operators) and
$Z_{i}=P_{i}(Z)=P_{i}\left(\sum_{k=1}^{\infty}\xi_{k}\mathfrak{h}_{k}\right)=P_{i}\left(\sum_{k=1}^{\infty}\xi_{k}i_{Z}\mathfrak{h}_{k}\right)=\sum_{k=1}^{\infty}\xi_{k}P_{i}(i_{Z}\mathfrak{h}_{k})=\sum_{k=1}^{\infty}\xi_{k}T_{i}(\mathfrak{h}_{k}),$
where we could apply $P_{i}$ term by term as the series converges in $L^{2}(\Omega;G)$. As $D\xi_{k}=DW(\mathfrak{h}_{k})=\mathfrak{h}_{k}$, we obtain
$DZ_{i}=\sum_{k=1}^{\infty}T_{i}(\mathfrak{h}_{k})\otimes\mathfrak{h}_{k}=T_{i}.$
Using $L^{-1}Z_{i}=Z_{i}$, we can hence compute
$\Gamma(Z_{i},-L^{-1}Z_{j})=\Gamma(Z_{i},Z_{j})=\widehat{C}_{\mathfrak{H}}(DZ_{i},DZ_{j})=\widehat{C}_{\mathfrak{H}}(T_{i},T_{j})=\sum_{k=1}^{\infty}\left(T_{j}(\mathfrak{h}_{k}),\cdot\right)_{E_{j},E_{j}^{\ast}}T_{i}(\mathfrak{h}_{k}),$
On the other hand, using the series representation (22), we obtain the same result:
$R_{Z_{i},Z_{j}}=\mathbb{E}\left(\left(Z_{j},\cdot\right)_{E_{j},E_{j}^{\ast}}Z_{i}\right)\\ =\mathbb{E}\left(\left(P_{j}Z,\cdot\right)_{E_{j},E_{j}^{\ast}}P_{i}Z\right)=\mathbb{E}\left(\left(\sum_{k=1}^{\infty}\xi_{k}T_{j}(\mathfrak{h}_{k}),\cdot\right)_{E_{j},E_{j}^{\ast}}\sum_{l=1}^{\infty}\xi_{l}T_{i}(\mathfrak{h}_{l}),\right)\\ =\sum_{k,l=1}^{\infty}\mathbb{E}\left(\xi_{k}\xi_{l}\left(T_{j}(\mathfrak{h}_{k}),\cdot\right)_{E_{j},E_{j}^{\ast}}T_{i}(\mathfrak{h}_{l})\right)=\sum_{k=1}^{\infty}\left(T_{j}(\mathfrak{h}_{k}),\cdot\right)_{E_{j},E_{j}^{\ast}}T_{i}(\mathfrak{h}_{k}).$
Analogous calculations (applying the tensor-valued definition of $C_{\mathfrak{H}}$) yields the identities for the $Q_{Z_{i},Z_{j}}$. ∎
Combining Propositions 3.6 and 3.7 yields a version of the well-known Gaussian integration by parts formula.
Corollary 3.8 (Gaussian integration by parts).
Let $X$ and $Y$ be jointly Radon Gaussian, taking values in Banach spaces $E$ and $F$, respectively, so that $(X,Y)$ is Radon and $f\in\mathcal{C}_{b}^{1}(E,F^{\ast})$. Then,
$\operatorname{tr}_{\pi}\left(Q_{f(X),Y}\right)=\left(\mathbb{E}\left(f^{\prime}(X)\right),Q_{X,Y}\right)_{\pi},$
where $f^{\prime}(x)$ is identified with an element of $\mathcal{L}(E,F;\mathbb{R})$ and $\left(\cdot,\cdot\right)_{\pi}$ is the pairing between the projective tensor product and its dual. If $F$ has the approximation property, we also have that
$\operatorname{tr}\left(R_{f(X),Y}\right)=\operatorname{tr}\left(\mathbb{E}\left(f^{\prime}(X)\right)R_{X,Y}\right).$
Remark 3.9.
Taking $E=F^{\ast}$ and $f=\operatorname{Id}_{F^{\ast}}$ in Proposition 3.6 yields that for $X\in\mathbb{D}^{1,2}(F^{\ast})$ and $Y\in\mathbb{D}^{1,2}(F)$ one has
$\operatorname{tr}_{\pi}\left(Q_{X,Y}\right)$ $=\mathbb{E}\left(\left(\operatorname{Id}_{F^{\ast}},\Gamma_{\pi}(X,-L^{-1}Y)\right)_{\pi}\right)$
$=\mathbb{E}\left(\operatorname{tr}_{\pi}\left(\Gamma_{\pi}\left(X,-L^{-1}Y\right)\right)\right)$
$=\mathbb{E}\left(\left(DX,-DL^{-1}Y\right)_{F^{\ast},F}\right).$
The pairing inside the last expectation is the trace duality pairing between the random $\gamma$-radonifying operators $DX\in\gamma(\mathfrak{H},F^{\ast})$ and $DL^{-1}Y\in\gamma(\mathfrak{H},F)$. Applying Lemma 3.1 then gives
$\mathbb{E}\left(\left(DX,-DL^{-1}Y\right)_{F^{\ast},F}\right)\leq\mathbb{E}\left(\left\lVert DX\right\rVert_{\gamma(\mathfrak{H},F^{\ast})}\left\lVert-DL^{-1}Y\right\rVert_{\gamma(\mathfrak{H},F)}\right),$
and if we take $X$ and $Y$ to be Gaussian, then by Proposition 3.7 the Malliavin derivatives become deterministic and the pseudo inverse of the generator disappears. As any two $\gamma$-radonifying operators $S\in\gamma(\mathfrak{H},F^{\ast})$ and $T\in\gamma(\mathfrak{H},F)$ can be written as the Malliavin derivatives of two Radon Gaussians $X$ and $Y$ such that $(X,Y)$ is jointly Gaussian (see the proof of Proposition 3.7), we recover the well-known continuity proof of the trace duality pairing for $\gamma$-radonifying operators (see Theorem 10.2 in  [vN10]).
On another note, taking $E=F$ and $f\in\mathcal{C}_{b}(E,E^{\ast})$, then by the above discussion, Lemma 2.5 and the ideal property of the $\gamma$-radonifying operators, we obtain
$\operatorname{tr}_{\pi}\left(Q_{f(X),X}\right)$ $\leq\mathbb{E}\left(\left\lVert Df(X)\right\rVert_{\gamma(\mathfrak{H},E^{\ast})}\left\lVert-DL^{-1}X\right\rVert_{\gamma(\mathfrak{H},E)}\right)$
$\leq\sqrt{\mathbb{E}\left(\left\lVert Df(X)\right\rVert_{\gamma(\mathfrak{H},E^{\ast})}^{2}\right)\mathbb{E}\left(\left\lVert-DL^{-1}X\right\rVert_{\gamma(\mathfrak{H},E)}^{2}\right)}$
$\leq\sqrt{\left\lVert f\right\rVert_{\mathcal{L}(E,E^{\ast})}\mathbb{E}\left(\left\lVert DX\right\rVert_{\gamma(\mathfrak{H},E)}^{2}\right)^{2}}$
$=\sqrt{\left\lVert f\right\rVert_{\mathcal{L}(E,E^{\ast})}}\mathbb{E}\left(\left\lVert DX\right\rVert_{\gamma(\mathfrak{H},E)}^{2}\right),$
which can be seen as some sort of generalized Poincaré inequality. In particular, taking $E=K$, where $K$ is some Hilbert space, $f$ the Riesz isometry and identifying $K$ with its dual gives
$\operatorname{tr}_{K}\left(R_{X}\right)$ $=\mathbb{E}\left(\left\lVert X-\mathbb{E}\left(X\right)\right\rVert_{K}^{2}\right)$
$\leq\mathbb{E}\left(\left\lVert DX\right\rVert_{\mathfrak{H}\widehat{\otimes}_{\sigma}K}\left\lVert-DL^{-1}X\right\rVert_{\mathfrak{H}\widehat{\otimes}_{\sigma}K}\right)$
$\leq\sqrt{\mathbb{E}\left(\left\lVert DX\right\rVert_{\mathfrak{H}\widehat{\otimes}_{\sigma}K}^{2}\right)\mathbb{E}\left(\left\lVert-DL^{-1}X\right\rVert_{\mathfrak{H}\widehat{\otimes}_{\sigma}K}^{2}\right)}$
$\leq\mathbb{E}\left(\left\lVert DX\right\rVert_{\mathfrak{H}\widehat{\otimes}_{\sigma}K}^{2}\right),$
which is the well-known vector valued Poincaré inequality (see for example  [Bog98, Rmk. 5.5.6]). Recall that $\mathfrak{H}\widehat{\otimes}_{\sigma}K$ can be identified with the Hilbert-Schmidt operators.
Now that an integration by parts formula is available, one can apply the well-known smart path method. Recall that the carré du champ operator can be defined as a random tensor-valued mapping or a random nuclear operator. We use the same symbol to denote both versions, and it will be clear from the context whether a tensor or operator is meant.
Proposition 3.10.
For a Banach space $E$, let $X\in\mathbb{D}^{1,2}(E)$ be centered and $Z$ be an independent, centered and $E$-valued Radon Gaussian random element with covariance tensor $Q_{Z}$ and covariance operator $R_{Z}$. Define, for $t\in[0,1]$,
$U_{t}=\sqrt{t}X+\sqrt{1-t}Z.$
Then, for all $f\in C^{2}_{b}(E)$, one has
(23) $\mathbb{E}(f(X))-\mathbb{E}(f(Z))=\frac{1}{2}\int_{0}^{1}\mathbb{E}\left(\left(f^{\prime\prime}(U_{t}),\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right)_{\pi}\right)dt$
and consequently
(24) $\left|\mathbb{E}(f(X))-\mathbb{E}(f(Z))\right|\leq\frac{1}{2}\left\lVert f^{\prime\prime}\right\rVert_{\infty}\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{\pi}.$
If $E$ has the approximation property, one also has
$\mathbb{E}(f(X))-\mathbb{E}(f(Z))=\frac{1}{2}\int_{0}^{1}\mathbb{E}\left(\operatorname{tr}\left(f^{\prime\prime}(U_{t})\left(\Gamma(X,-L^{-1}X)-R_{Z}\right)\right)\right)dt$
and
$\left|\mathbb{E}(f(X))-\mathbb{E}(f(Z))\right|\leq\frac{1}{2}\left\lVert f^{\prime\prime}\right\rVert_{\infty}\left\lVert\Gamma(X,-L^{-1}X)-R_{Z}\right\rVert_{\nu}.$
Proof.
We prove the identities for $\Gamma_{\pi}$. The corresponding versions for $\Gamma$ follow with the usual identifications. Define $\Phi(t)=\mathbb{E}\left(f(U_{t})\right)$, so that
$\mathbb{E}\left(f(X)\right)-\mathbb{E}\left(f(Z)\right)=\int_{0}^{1}\Phi^{\prime}(t)dt.$
Because $f$ and its derivative is bounded and $X,Z\in L^{2}(\Omega;E)$, we may differentiate inside of the expectation and obtain
$\Phi^{\prime}(t)$ $=\mathbb{E}\left(f^{\prime}(U_{t})\left(\frac{X}{2\sqrt{t}}-\frac{Z}{2\sqrt{1-t}}\right)\right)$
$=\frac{1}{2\sqrt{t}}\mathbb{E}\left((X,f^{\prime}(U_{t}))_{E,E^{\ast}}\right)-\frac{1}{2\sqrt{1-t}}\mathbb{E}\left((Z,f^{\prime}(U_{t}))_{E,E^{\ast}}\right)$
$=\frac{1}{2\sqrt{t}}\mathbb{E}\left(\operatorname{tr}_{\pi}\left(Q_{f^{\prime}(\sqrt{t}X+z),X}\right)_{\mid z=\sqrt{1-t}Z}\right)$
(25) $\qquad\qquad-\frac{1}{2\sqrt{1-t}}\mathbb{E}\left(\operatorname{tr}_{\pi}\left(Q_{f^{\prime}(x+\sqrt{1-t}Z),Z}\right)_{\mid x=\sqrt{t}X}\right)$
Integrating both terms by parts, i.e. applying Proposition 3.6, yields
$\frac{1}{2\sqrt{t}}\mathbb{E}\left(\operatorname{tr}_{\pi}\left(Q_{f^{\prime}(\sqrt{t}X+z),X}\right)_{\mid z=\sqrt{1-t}Z}\right)=\sqrt{t}\,\mathbb{E}\left(\left(f^{\prime\prime}(U_{t}),\Gamma_{\pi}(X,-L^{-1}X)\right)_{\pi}\right)$
and
$\frac{1}{2\sqrt{1-t}}\mathbb{E}\left(\operatorname{tr}_{\pi}\left(Q_{f^{\prime}(x+\sqrt{1-t}Z),Z}\right)_{\mid x=\sqrt{t}X}\right)=\sqrt{1-t}\,\mathbb{E}\left(\left(f^{\prime\prime}(U_{t}),Q_{Z}\right)_{\pi}\right).$
Plugged into (25), we obtain
$\Phi^{\prime}(t)=\frac{1}{2}\mathbb{E}\left(\left(f^{\prime\prime}(U_{t}),\Gamma_{\pi}(X,-L_{E}^{-1}X)-Q_{Z}\right)_{\pi}\right),$
which proves (23). Furthermore,
$\left|\left(f^{\prime\prime}(U_{t}),\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right)_{\pi}\right|$ $\leq\left\lVert f^{\prime\prime}(U_{t})\right\rVert_{\mathcal{L}(E,E;\mathbb{R})}\,\left\lVert\Gamma_{\pi}(X,-L^{-1}_{E}X)-Q_{Z}\right\rVert_{\pi}$
$\leq\left\lVert f^{\prime\prime}\right\rVert_{\infty}\,\left\lVert\Gamma_{\pi}(X,-L^{-1}_{E}X)-Q_{Z}\right\rVert_{\pi}.$
Together with (23), this yields (24). ∎
4. Bounds in the bounded Lipschitz metric via regularization
In this section, we combine the bound (24) obtained via integration by parts with the regularization procedure developed in [BRZ24] in order to derive bounds for the bounded Lipschitz metric on the Skorokhod space (or any closed subspace thereof). The regularization procedure is designed to establish a link between this metric and a smooth integral probability metric initially introduced in [Bar90], which uses Fréchet differentiable functions and is therefore in general only able to guarantee convergence of finite-dimensional distributions. This smooth distance can be controlled by the difference between the projective norm of the carré du champ and the Gaussian covariance tensor via (24). The regularization introduces additional terms ensuring tightness in such a way, that under mild assumptions global control by the aforementioned projective norm can be achieved via parameter optimization.
Recall that the bounded Lipschitz metric $d_{BL}$ on the class of Borel probability measures on a Banach space $E$ is defined as
$d_{BL}(\mu,\nu)=\sup_{f\in\operatorname{BL}(E)}\left|\int_{E}f(x)\mu(dx)-\int_{E}f(x)\nu(dx)\right|,$
where $\operatorname{BL}(E)$ denotes the set of all bounded Lipschitz functions $f\colon E\to\mathbb{R}$ with Lipschitz constant at most one, such that $\left\lVert f\right\rVert_{\infty}\leq 1$.
For two $E$-valued Borel random elements, we define $d_{BL}(X,Y)=d_{BL}(P_{X},P_{Y})$, where $P_{X}$ and $P_{Y}$ denote the laws of $X$ and $Y$.
Throughout this section, we fix $T>0$, $d\in\mathbb{N}$ and denote by $\mathcal{D}$ the Skorokhod space $D([0,T];\mathbb{R}^{d})$ equipped with the uniform metric.
As in [Bar90], we define $M^{0}\subseteq\mathcal{C}_{b}^{2}(\mathcal{D})$ to consist of all twice continuously Fréchet differentiable functions $f\colon\mathcal{D}\to\mathbb{R}$ such that $f$ and its first two derivatives are uniformly bounded and $f^{\prime\prime}$ is Lipschitz continuous, i.e. for which the norm $\left\lVert f\right\rVert_{M^{0}}$ defined by
$\left\lVert f\right\rVert_{M^{0}}=\left\lVert f\right\rVert_{\mathcal{C}_{b}^{2}}+\left\lVert f^{\prime\prime}\right\rVert_{\text{Lip}}=\left\lVert f\right\rVert_{\infty}+\left\lVert f^{\prime}\right\rVert_{\infty}+\left\lVert f^{\prime\prime}\right\rVert_{\infty}+\left\lVert f^{\prime\prime}\right\rVert_{\text{Lip}}$
is finite.
The $\varepsilon$-regularization introduced in [BRZ24], is defined as follows.
Definition 4.1.
For $\varepsilon>0$, a Banach space $E$ and a Bochner integrable function $f\colon[0,T]\to E$, we define its $\varepsilon$-regularized version $f_{\varepsilon}$ by
$f_{\varepsilon}(x)=\frac{1}{2\varepsilon}\int_{x-\varepsilon}^{x+\varepsilon}f\left([x]_{[0,T]}\right)\,dx,$
where $\cdot]_{[0,T]}=\operatorname{min}\left\{\operatorname{max}\left\{0,\cdot\right\},T\right\}$ is the truncation operator, forcing its argument to stay in the interval $[0,T]$.
The following lemma is straightforward to see.
Lemma 4.2.
For a Banach space $E$, $p\geq 1$ and $T>0$, let $X=(X(t))_{t\in[0,T]}$ be a jointly measurable stochastic process such that $X(t)\in L^{p}(\Omega;E)$ for all $t\in[0,T]$. Then the following is true:
  1. (i)
    If  $\lim_{\varepsilon\to 0}\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}^{p}\right)=0$, then $X$ has almost surely continuous sample paths.
  2. (ii)
    Conversely, if $X$ has almost surely continuous sample paths and $\mathbb{E}\left(\left\lVert X\right\rVert_{\infty}^{p}\right)<\infty$, then $\lim_{\varepsilon\to 0}\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}^{p}\right)=0$.
Now we can formulate the mentioned link between the smooth integral probability metric based on $M^{0}$ and the bounded Lipschitz distance on $\mathcal{D}$. The following theorem is a special case of [BRZ24, Thm. 1.1 and Cor. 1.3] adjusted for our purposes.
Theorem 4.3.
Let $X$ and $Z$ be $\mathcal{D}$-valued random elements such that $Z$ has almost surely continuous sample paths. For $\varepsilon>0$, let $X_{\varepsilon}$ and $Z_{\varepsilon}$ be their $\varepsilon$-regularizations and suppose there exists $\kappa>0$ such that, for any $h\in M^{0}$, we have
(26) $\left|\mathbb{E}(h(X))-\mathbb{E}(h(Z))\right\rvert\leq\kappa\left\lVert h\right\rVert_{M^{0}}.$
Then, for any $\varepsilon>0$ and $\delta>0$, one has
$d_{BL}\left(X,Z\right)\leq\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}\right)+\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)+2\left(\sqrt{T}s_{d}+\sqrt{d}\right)\delta+\frac{4(T+2)}{\varepsilon^{2}\delta^{2}}\,\kappa,$
where $s_{d}=\mathbb{E}\left(\sup_{t\in[0,1]}\left\lVert B_{t}\right\rVert_{\mathbb{R}^{d}}\right)$ is the supremum of the norm of a $d$-dimensional standard Brownian motion.
In the above special case, one can immediately optimize in $\delta$ and then use our bound from Proposition 3.10, which satisfies (26) with
$\kappa=\frac{1}{2}\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{L^{1}(\Omega;\mathcal{D})}$
as $\left\lVert h^{\prime\prime}\right\rVert_{\infty}\leq\left\lVert h\right\rVert_{M^{0}}$. This yields the following result.
Theorem 4.4.
Let $X$ and $Z$ be centered $\mathcal{D}$-valued Radon random elements such that $Z$ is Radon Gaussian with almost surely continuous sample paths and $X\in\mathbb{D}^{1,2}(\mathcal{D})$. Then, for any $\varepsilon>0$, one has
$d_{BL}(X,Z)\leq\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}\right)+\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)+\frac{C_{T,d}}{\varepsilon^{2/3}}\left\lVert\Gamma_{\pi}(X,-L_{E}^{-1}X)-Q_{Z}\right\rVert_{L^{1}(\Omega;\mathcal{D}\widehat{\otimes}_{\pi}\mathcal{D})}^{1/3},$
where
$C_{T,d}=\frac{3}{2}\sqrt[3]{4(T+2)\left(\sqrt{T}s_{d}+\sqrt{d}\right)}.$
Remark 4.5.
As already mentioned, Theorem 4.3 is a special case of [BRZ24, Thm. 1.1 and Cor. 1.3], where in particular a bound for the Lévy-Prokhorov metric $d_{LP}$ is derived as well. In principle, we could apply the same strategy as for the bounded Lipschitz metric, but as the parameter optimization is more involved without concrete estimates for the regularization terms and in general (see  [Dud68]) one has that
(27) $\frac{2}{3}d_{LP}\leq d_{BL}\leq 2d_{LP},$
we refrained from doing so.
Remark 4.6.
In the setting of Theorem 4.4, if we assume that there exists a non-negative function $\varphi\in C([0,\infty))$ vanishing at zero such that
$\operatorname{max}\left\{\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}\right),\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)\right\}\leq\varphi(\varepsilon)$
for all $\varepsilon>0$, this implies the existence of a continuous function $\psi$ vanishing at zero, such that
(28) $d_{BL}(X,Z)\leq\psi\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{L^{1}(\Omega;\mathcal{D}\widehat{\otimes}_{\pi}\mathcal{D})}\right).$
Analogously, taking a sequence $(X_{n})$ instead of $X$ and assuming
(29) $\operatorname{max}\left\{\mathbb{E}\left(\left\lVert X_{n,\varepsilon}-X_{n}\right\rVert_{\infty}\right),\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)\right\}\leq\varphi(\varepsilon)+a_{n}$
for all $\varepsilon>0$ and $n\in\mathbb{N}$ such that $a_{n}\geq 0$ and $\lim_{n\to\infty}a_{n}=0$ yields the existence of a continuous function $\psi$ vanishing in zero satisfying
$d_{BL}(X_{n},Z)\leq a_{n}+\psi\left(\left\lVert\Gamma_{\pi}(X_{n},-L_{E}^{-1}X_{n})-Q_{Z}\right\rVert_{L^{1}(\Omega;\mathcal{D}\widehat{\otimes}_{\pi}\mathcal{D})}\right),$
In both scenarios, this can be seen by choosing $\varepsilon=\kappa^{\alpha}$ for some $\alpha\in(0,1/2)$. A similar statement is true in the more general case when the left hand side of (29) is bounded by suitable functions $\varphi_{n}(\varepsilon)$.
As a somewhat more concrete example, we carry out the $\varepsilon$-optimization for the case where the $\varepsilon$-regularizations have a Hölder-continuous modulus of continuity.
Proposition 4.7.
In the setting of Theorem 4.4, assume there exist positive constants $M$ and $\alpha$ such that one has
$\operatorname{max}\left\{\mathbb{E}\left(\left\lVert X_{\varepsilon}-X\right\rVert_{\infty}\right),\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)\right\}\leq M\varepsilon^{\alpha}$
for all $\varepsilon>0$. Then
$d_{BW}\left(X,Z\right)\leq C_{\alpha,T,M,d}\,\mathbb{E}\left(\left\lVert\Gamma_{\pi}(X,-L^{-1}X)-Q_{Z}\right\rVert_{\pi}\right)^{\frac{\alpha}{\alpha+2/3}},$
where
$C_{\alpha,T,M,d}=\frac{M(2+3\alpha)}{2(M\alpha)^{\frac{3\alpha}{3\alpha+2}}}\left(4(T+2)\left(\sqrt{T}s_{d}+\sqrt{d}\right)\right)^{\frac{\alpha}{\alpha+2/3}}.$
Of course, if the joint topological support $E$ of the Radon random elements $X$ and $Z$ is a closed subspace of $\mathcal{D}$, most notably the space $\mathcal{C}([0,T];\mathbb{R}^{d})$ of continuous functions, all our bounds for $d_{BL}(X,Z)$ and $d_{LP}(X,Z)$ (via (27)) remain valid and cannot be improved by our approach when restricting both metrics to $E$. For $d_{LP}$ this is clear, as in the defining infimum only Borel sets contained in $E$ have to be considered. For $d_{BL}$, not that on the one hand, for any $f\in BL(\mathcal{D})$ its restriction $f_{\restriction E}$ is in $BL(E)$, so that $d_{\operatorname{BL}(\mathcal{D})}(X,Z)\leq d_{\operatorname{BL}(E)}(X,Z)$. On the other hand, for $g\in BL(E)$, its McShane-Whitney extension $\widetilde{g}$ (see [McS34, Theorem 1] or [Whi34]) to $\mathcal{D}$ yields a function in $BL(\mathcal{D})$ as it preserves the values of both, the Lipschitz- and the uniform norm. This shows $d_{BL(E)}(X,Z)\leq d_{BL(\mathcal{D})}(X,Z)$.
Finally, we recall the well-known facts that the uniform topology on $\mathcal{D}$ is stronger than all of the four standard Skorokhod topologies, and that uniform convergence in $C([0,T];\mathbb{R}^{d})$ is equivalent to convergence in the $J^{1}$-topology.
5. Contraction bounds on Wiener chaos
In this section, we show how the carré du champ norms appearing in the bounds in Section 4 can be controlled in terms of kernel contractions, when the random element $X$ is an element of a Wiener chaos (or a finite sum thereof). Throughout, $E$ and $F$ will be real Banach spaces and $\mathfrak{H}$ a real, separable Hilbert space.
Recall from Section 2.5 the space $\gamma^{p}(\mathfrak{H},E)$ of iterated $\gamma$-radonifying $E$-valued mappings and the subspace $\gamma_{sym}^{p}(\mathfrak{H},E)$ of its symmetrization, and from Section 2.7 the definition of the multiple integral $I_{p}(f)$ for $f\in\gamma_{sym}^{p}(\mathfrak{H},E)$.
We will show that, as in finite-dimension, the carré du champ of a multiple integral $I_{p}(f)$ decomposes into a finite sum of multiple integrals whose kernels are given by contractions of $f$. This will then allow to control the nuclear norm by these contractions.
For kernels $f\in\gamma^{p}(\mathfrak{H},E)$ and $g\in\gamma^{p}(\mathfrak{H},E)$, we denote the symmetrizations of the tensor- and operator-valued contractions (see Lemma 3.1 for their definition) by
$\widehat{C}_{\mathfrak{H},sym}^{r}(f,g)=\operatorname{Sym}\widehat{C}_{\mathfrak{H}}(f,g)$
and
$\widehat{C}_{\mathfrak{H},sym,\pi}^{r}(f,g)=\operatorname{Sym}\widehat{C}_{\mathfrak{H},\pi}(f,g)$
respectively, where $\operatorname{Sym}$ is the symmetrization operator acting on the q+p2rq+p-2r remaining $\mathfrak{H}$-variables (see Subsection 2.5).
As in the scalar-case, the linear symmetrization operator is continuous, i.e. the norm of a symmetrized contraction is bounded by its non-symmetrized norm (see [Maa10, Thm. 4.2]), so that together with Lemma 3.1 we have
$\left\lVert\widehat{C}_{\mathfrak{H},\pi,sym}^{r}(f,g)\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$ $\leq\left\lVert\widehat{C}_{\mathfrak{H},\pi}^{r}(f,g)\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$
$\leq\left\lVert f\right\rVert_{\gamma^{p}(\mathfrak{H},E)}\left\lVert g\right\rVert_{\gamma^{q}(\mathfrak{H},F)}.$
and the same inequalities hold for the operator-valued counterpart.
When no ambiguity can arise, we will follow the de facto standard notational convention of the Malliavin community and write $f\otimes_{r}g$ for $\widehat{C}_{\mathfrak{H}}^{r}(f,g)$ and $f\widetilde{\otimes}_{r}g$ for its symmetrization. The tensor valued counterparts will be denoted by $f\otimes_{r,\pi}g$ and $f\widetilde{\otimes}_{r,\pi}g$, respectively. Note that the familiar scalar-valued contractions are obtained by choosing $E=F=\mathbb{R}$, as $\mathbb{R}\widehat{\otimes}_{\pi}\mathbb{R}\simeq\mathfrak{N}(\mathbb{R},\mathbb{R})\simeq\mathbb{R}$ and $\gamma^{p}(\mathfrak{H},\mathbb{R})\simeq\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}$, and analogously $\gamma^{p}_{sym}(\mathfrak{H},\mathbb{R})\simeq\mathfrak{H}^{\widehat{\odot}_{\sigma}p}$
As a caveat, we would like to mention again that for $\mathfrak{H}=L^{2}(M)=L^{2}(M,\mathcal{M},\mu)$, with $\mu$ a $\sigma$-finite measure without atoms on a measurable space $(M,\mathcal{M})$, one has to work with representable operators (see Section 2.7). For the corresponding representing functions however, the contraction is obtained in the same way as in the scalar-valued case, i.e. by identifying and then integrating out variables.
We now extend the product formula (see for example [Nua06, Prop. 1.1.3] or  [NP12, Thm. 2.7.10]) to the vector-valued setting. As in the previous sections, given $u\in E$ and $v\in F$, we denote by $u\otimes v$ both the simple tensor in $E\otimes F$ and the rank one operator in $\mathcal{F}(G^{\ast},E)$. Which definition is meant can always be inferred from the context.
Proposition 5.1.
For $m\in(1,\infty)$, $p,q\in\mathbb{N}$, $f\in\gamma^{p}_{sym}(\mathfrak{H},E)$ and $g\in\gamma^{q}_{sym}(\mathfrak{H},F)$, one has
(30) $I_{p}(f)\otimes I_{q}(g)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right)$
in $L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)$ and, interpreting the left hand side as a random rank-one operator,
$I_{p}(f)\otimes I_{q}(g)=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r}g\right)$
in $L^{m}(\Omega;\mathfrak{N}(F^{\ast},E))$.
Proof.
We first show the claim for the tensor-valued contractions. Let $\left\{\mathfrak{h}_{k}\colon k\in\mathbb{N}\right\}$ be an orthonormal basis of $\mathfrak{H}$. For finite rank kernels $f\in\mathcal{F}_{sym}^{p}(\mathfrak{H},E)\simeq\mathfrak{H}^{\odot p}\otimes E$ and $g\in\mathcal{F}_{sym}^{q}(\mathfrak{H},F)\simeq\mathfrak{H}^{\odot q}\otimes F$, we have representations as in (9) of the form
$f=\sum_{\mathbf{j}\in[n]^{p}}\mathfrak{h}_{\mathbf{j}}\otimes x_{\mathbf{j}}\qquad\text{and}\qquad g=\sum_{\mathbf{k}\in[n]^{q}}\mathfrak{h}_{\mathbf{k}}\otimes y_{\mathbf{k}},$
where $n\in\mathbb{N}$, $x_{\mathbf{j}}\in E$ and $y_{\mathbf{k}}\in F$ for $\mathbf{j}\in[n]^{p}$ and $\mathbf{k}\in[n]^{q}$, respectively. For clarity, and just in this proof, we will add the codomain of the kernel the operator $I_{p}$ acts on as a superscript, writing for example $I_{p}^{E}(f)$ instead of $I_{p}(f)$. Analogously, we add the codomains of the kernels to the contraction operation, writing for example $f\widetilde{\otimes}_{r,\pi}^{E,F}g$ instead of $f\widetilde{\otimes}_{r,\pi}g$.
Then, by the product formula for real-valued kernels (see for example  [Nua06, Prop. 1.1.3] or [NP12, Thm. 2.7.10]),
$I_{p}^{E}(f)\otimes I_{q}^{F}(g)$ $=\sum_{\mathbf{j}\in[n]^{p},\mathbf{k}\in[n]^{q}}I_{p}^{\mathbb{R}}(\mathfrak{h}_{\mathbf{j}})I_{q}^{\mathbb{R}}(\mathfrak{h}_{\mathbf{k}})x_{\mathbf{j}}\otimes y_{\mathbf{k}}$
$=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}\sum_{\mathbf{j}\in[n]^{p},\mathbf{k}\in[n]^{q}}I_{p+q-2r}^{\mathbb{R}}\left(\mathfrak{h}_{\mathbf{j}}\widetilde{\otimes}_{r}^{\mathbb{R},\mathbb{R}}\mathfrak{h}_{\mathbf{k}}\right)x_{\mathbf{j}}\otimes y_{\mathbf{k}}$
$=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}^{E\widehat{\otimes}_{\pi}F}\left(\sum_{\mathbf{j}\in[n]^{p},\mathbf{k}\in[n]^{q}}\left(\mathfrak{h}_{\mathbf{j}}\widetilde{\otimes}_{r}^{\mathbb{R},\mathbb{R}}\mathfrak{h}_{\mathbf{k}}\right)\,x_{\mathbf{j}}\otimes y_{\mathbf{k}}\right)$
$=\sum_{r=0}^{p\land q}r!\binom{p}{r}\binom{q}{r}I_{p+q-2r}^{E\widehat{\otimes}_{\pi}F}\left(f\widetilde{\otimes}_{r,\pi}^{E,F}g\right).$
For general kernels $f\in\gamma^{p}(\mathfrak{H},E)$ and $g\in\gamma^{q}(\mathfrak{H},F)$, we approximate by finite-rank sequences $(f_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}^{p}_{sym}(\mathfrak{H},E)$ and $(g_{n})_{n\in\mathbb{N}}\subseteq\mathcal{F}^{q}_{sym}(\mathfrak{H},F)$. The Wiener-Itô norm equivalence (see (14) in Section 2.7) and the fact that changing the $L^{2}$-norm in the $\gamma$-radonifying norm to an $L^{m}$ norm for some $m\in[1,\infty)$ yields an equivalent $\gamma$-norm due to the Kahane-Khintchine inequalities, we obtain that $I_{p}(f_{n})\to I_{p}(f)$ in $L^{m}(\Omega;E)$ and also that $I_{q}(g_{n})\to I_{q}(g)$ in $L^{m}(\Omega;F)$ for $m\in[1,\infty)$. Together with boundedness of the contraction operator (see (17)) and of the multiple integral operator $I_{p}$, this also yields that
$I_{p+q-2r}^{E\widehat{\otimes}_{\pi}F}\left(f_{n}\widetilde{\otimes}_{r,\pi}^{E,F}g_{n}\right)\to I_{p+q-2r}^{E\widehat{\otimes}_{\pi}F}\left(f\widetilde{\otimes}_{r,\pi}^{E,F}g\right)$
in $L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)$. ∎
Remark 5.2.
By the universal property of the projective tensor product, the “multiplication map”
$T\colon\mathcal{H}_{p}(E)\otimes\mathcal{H}_{q}(F)\to L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)$
induced by (30) extends isometrically to $\mathcal{H}_{p}(E)\widehat{\otimes}_{\pi}\mathcal{H}_{q}(F)$.
Via the product formula, one can now obtain a decomposition of the carré du champ into a sum of multiple integrals.
Proposition 5.3.
For $p,q\in\mathbb{N}$, $f\in\gamma^{p}_{sym}(\mathfrak{H},E)$ and $g\in\gamma^{q}_{sym}(\mathfrak{H},F)$, one has
$\Gamma_{\pi}\left(I_{p}(f),I_{q}(g)\right)=\sum_{r=1}^{p\land q}a_{p,q,r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right)$
and
$\Gamma\left(I_{p}(f),I_{q}(g)\right)=\sum_{r=1}^{p\land q}a_{p,q,r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r}g\right),$
where
(31) $a_{p,q,r}=pq(r-1)!\binom{p-1}{r-1}\binom{q-1}{r-1}$.
Proof.
By the same density argument as in the proof of Proposition 5.1, it suffices to show the claim for $f\in\mathfrak{H}^{\odot p}\otimes E$ and $g\in\mathfrak{H}^{\odot q}\otimes F$, so that
$I_{p}(f)=\sum_{j=1}^{n}I_{p}(f_{j})\otimes x_{j}\qquad\text{and}\qquad I_{q}(g)=\sum_{k=1}^{n}I_{q}(g_{k})\otimes y_{k}$
with $f_{j}\in\mathfrak{H}^{\odot p}$, $g_{k}\in\mathfrak{H}^{\odot q}$, $x_{j}\in E$ and $y_{k}\in F$ for $1\leq j,k\leq n$. Then, by bilinearity,
$\Gamma_{\pi}(I_{p}(f),I_{q}(g))=\sum_{j,k=1}^{n}\Gamma\left(I_{p}(f_{j}),I_{q}(g_{k})\right)x_{j}\otimes y_{k}.$
The claim now follows from the scalar-valued identity
$\Gamma\left(I_{p}(f_{j}),I_{q}(g_{k})\right)=\sum_{r=1}^{p\land q}a_{p,q,r}I_{p+q-2r}\left(f_{j}\widetilde{\otimes}_{r}g_{k}\right),$
which is a direct consequence of the product formula (see for example [NP09, Prop. 3.2], [NPR10, Lem. 3.7] or [NP12, Proof of Lem. 6.2.1]). Identifying $x_{j}\otimes y_{k}$ with a rank-one operator yields the operator valued counterpart. ∎
We are now in a position to prove the main result of this section, namely a bound for the nuclear norm of the random covariance difference in terms of norms of contracted kernels. Note that, up to the constant $c_{p,q,m}$, these contraction bounds reduce to the classical scalar-valued bounds for $E=F=\mathbb{R}$ (which can be found in the references given above in the proof of Proposition 5.1).
Theorem 5.4.
For $m\in[1,\infty)$, $p,q\in\mathbb{N}$, $f\in\gamma^{p}_{sym}(\mathfrak{H},E)$ and $g\in\gamma^{q}_{sym}(\mathfrak{H},F)$, let $X=I_{p}(f)$, $Y=I_{q}(g)$ and denote the cross-covariance tensor and operator of $X$ and $Y$ by $Q_{X,Y}$ and $R_{X,Y}$, respectively. Then
$\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-Q_{X,Y}\right\rVert_{L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)},$
and
$\left\lVert\Gamma(X,-L^{-1}Y)-R_{X,Y}\right\rVert_{L^{2}(\Omega;\mathfrak{N}(F^{\ast},E))}\\ \leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r}g\right\rVert_{\gamma^{p+q-2r}(\mathfrak{H},\mathfrak{N}(F^{\ast},E))},$
where the positive constants $\mathfrak{C}_{m,2}$ and $a_{p,q,r}$ are defined at (14) and (31), respectively.
Proof.
By Proposition 5.3, one has
$\Gamma_{\pi}(X,-L^{-1}Y)=\frac{1}{q}\Gamma_{\pi}(X,Y)=\frac{1}{q}\sum_{r=1}^{p\land q}a_{p,q,r}\,I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right).$
If $p\neq q$, then both $Q_{X,Y}$ and $R_{X,Y}$ vanish as elements from different Wiener chaoses are uncorrelated (this transfers from the finite-dimensional case by approximation). If $p=q$, we can take expectation on both sides of the above sum and obtain by Lemma 3.5 and the fact that all multiple integrals of positive order are centered, that
$Q_{X,Y}=\mathbb{E}\left(\Gamma_{\pi}(X,-L^{-1}Y)\right)=\frac{a_{p,q,p}}{q}I_{0}\left(f\widetilde{\otimes}_{p,\pi}g\right).$
Therefore
$\Gamma_{\pi}(X,-L^{-1}Y)-Q_{X,Y}=\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right),$
so that by the triangle inequality and the Wiener-Itô equivalence of norms (see (14) in Section 2.7),
$\left\lVert\Gamma_{\pi}(X,-L^{-1}Y)-Q_{X,Y}\right\rVert_{L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)}$
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\left\lVert I_{p+q-2r}\left(f\widetilde{\otimes}_{r,\pi}g\right)\right\rVert_{L^{m}(\Omega;E\widehat{\otimes}_{\pi}F)}$
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\widetilde{\otimes}_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$
$\qquad\qquad\leq\frac{1}{q}\sum_{r=1}^{p\land q-1}a_{p,q,r}\,\mathfrak{C}_{p+q-2r,m}\,\left\lVert f\otimes_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}$
The operator-valued version follows analogously. ∎
Recall from Section 2.5 that the $E$-valued Wiener chaoses contain nuclear series. Given two such nuclear series $I_{p}(f)$ and $I_{q}(g)$ of the form
$I_{p}(f)=\sum_{j=1}^{\infty}I_{p}(f_{j})\otimes x_{j}\qquad\text{and}\qquad I_{q}(g)=\sum_{k=1}^{\infty}I_{q}(g_{k})\otimes y_{k}$
with $f_{j}\in\mathfrak{H}^{\widehat{\odot}_{\sigma}p}$, $g_{k}\in\mathfrak{H}^{\widehat{\odot}_{\sigma}q}$, $x_{j}\in E$ and $y_{k}\in F$ for all $j,k\in\mathbb{N}$, it is readily obtained that for $1\leq r\leq p\land q-1$ one has
$\left\lVert f\otimes_{r,\pi}g\right\rVert_{\gamma^{(p+q-2r)}(\mathfrak{H},E\widehat{\otimes}_{\pi}F)}^{2}\leq\sum_{j,k=1}^{\infty}\left\lVert f_{j}\otimes_{r}g_{k}\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p+q-2r}}\left\lVert x_{j}\right\rVert_{E}\left\lVert y_{k}\right\rVert_{F},$
where the right hand side is finite as
$\left\lVert f_{j}\otimes_{r}g_{k}\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p+q-2r}}\leq\left\lVert f_{j}\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}}\left\lVert g_{k}\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}q}}.$
In this special case, one can hence reduce the computation of vector valued contractions to the well-studied scalar case. This is in particular true for finite-rank kernels, in which case all above sums are finite.
To end this section, we illustrate using a simple example how the abstract contraction bounds can be combined with the results of the previous section to bounds the bounded Lipschitz distance between a multiple integral and a Gaussian taking values in the space of continuous functions.
To prepare this example, we first prove that Hölder-continuous paths provide control on the $\varepsilon$-regularization.
Proposition 5.5.
Let $W$ be an isonormal Gaussian process on a real, separable Hilbert space $\mathfrak{H}$. Fix $p\geq 1$, $T>0$ and let $f\in\gamma^{p}(\mathfrak{H},\mathcal{C}([0,T]))$ with Hölder continuous range, i.e. there exists $L>0$ and $\beta\in(0,1]$ such that for all $x,y\in[0,T]$ one has
$\left\lVert(e_{y}-e_{x})\circ f\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}}\leq L\left|y-x\right|^{\beta},$
where $e_{x}$ denotes the point evaluation functional at $x\in[0,T]$. Then, for all $s>\max\left\{2,\frac{1}{\beta}\right\}$, all $\eta\in(0,\beta-\frac{1}{s})$ and all $\varepsilon\in(0,T)$, one has
$\mathbb{E}\left(\left\lVert\left(I_{p}(f)\right)_{\varepsilon}-I_{p}(f)\right\rVert_{\infty}\right)\leq C\,L\,\varepsilon^{\beta-\frac{1}{s}-\eta},$
where $\left(I_{p}(f)\right)_{\varepsilon}$ is the $\varepsilon$-regularization of $I_{p}(f)$ (see Definition 4.1) and the positive constant $C$ is given by
$C=8\left(p-1\right)^{\frac{s}{2}}\sqrt{p!}\left(\frac{2}{\eta s}\right)^{\frac{1}{s}}T^{\frac{1+\eta s}{s}}.$
Remark 5.6.
Note that as $s$ can be chosen arbitrarly large and $\eta>0$ arbitrarly close to zero, we have that for every $\zeta\in(0,\beta)$, there exists $C_{\zeta}>0$ such that
$\mathbb{E}\left(\left\lVert\left(I_{p}(f)\right)_{\varepsilon}-I_{p}(f)\right\rVert\right)\leq C_{\zeta}\,L\,\varepsilon^{\beta-\zeta}.$
Proof.
Applying the Wiener-Itô isometry, we can write
$\left\lVert(e_{y}-e_{x})\circ I_{p}(f)\right\rVert_{L^{2}(\Omega)}=\left\lVert I_{p}\left((e_{y}-e_{x})\circ f\right)\right\rVert_{L^{2}(\Omega)}\\ =\sqrt{p!}\left\lVert\left(e_{y}-e_{x}\right)\circ f\right\rVert_{\mathfrak{H}^{\widehat{\otimes}_{\sigma}p}}\leq\sqrt{p!}L\left|y-x\right|^{\beta}.$
Together with the hypercontractivity property, this yields for any $s\geq 2$ that
(32) $\left\lVert(e_{y}-e_{x})\circ I_{p}(f)\right\rVert_{L^{s}(\Omega)}\leq\left(p-1\right)^{s/2}\left\lVert(e_{y}-e_{x})\circ I_{p}(f)\right\rVert_{L^{2}(\Omega)}\leq\left(p-1\right)^{s/2}\sqrt{p!}L\left|y-x\right|^{\beta}.$
Now, fix $x\in[0,T]$. By definition of the $\varepsilon$-regularization and monotonicity of expectation, we obtain
$\left|\left(I_{p}\left(e_{x}\circ f\right)\right)_{\varepsilon}-I_{p}\left(e_{x}\circ f\right)\right|$ $=\left|\mathbb{E}_{U}\left(I_{p}\left(\left(e_{[x+\varepsilon U]_{[0,T]}}-e_{x}\right)\circ f\right)\right)\right|$
$\leq\mathbb{E}_{U}\left(\left|I_{p}\left(\left(e_{[x+\varepsilon U]_{[0,T]}}-e_{x}\right)\circ f\right)\right|\right),$
where $U$ is uniformly distributed on (1,1)(-1,1), independent of $I_{p}(f))$, $\cdot]_{[0,T]}$ denotes the truncation operator and $\mathbb{E}_{U}$ is a shorthand for taking expectation with respect to $U$. Since $U\in(-1,1)$, one has $\left|x+\varepsilon U-x\right|\leq\varepsilon$. If $x+\varepsilon U\notin[0,T]$, it still follows that $\left|[x+\varepsilon U]_{[0,T]}-x\right|\leq\varepsilon$, since in this case $x+\varepsilon U]_{[0,T]}\in\left\{0,T\right\}$ and $x$ lies within $\varepsilon$-distance of the boundary. Therefore, there exists $h\in[-\varepsilon,\varepsilon]$ such that $x+h\in[0,T]$ and
$\left|I_{p}\left(\left(e_{[x+\varepsilon U]_{[0,T]}}-e_{x}\right)\circ f\right)\right|\leq\sup_{\left|h\right|\leq\varepsilon}\left|I_{p}\left(\left(e_{x+h}-e_{x}\right)\circ f\right)\right|$
Taking the supremum over $x$ and expectation, we get
(33) $\mathbb{E}\left(\left\lVert\left(I_{p}\left(e_{x}\circ f\right)\right)_{\varepsilon}-I_{p}\left(e_{x}\circ f\right)\right\rVert_{\infty}\right)\leq\mathbb{E}\left(\omega(I_{p}(f),\varepsilon)\right),$
where
$\omega\left(I_{p}(f),\varepsilon\right)=\sup_{\begin{subarray}{c}x,y\in[0,T]\\ \left|y-x\right\rvert\leq\varepsilon\end{subarray}}\left|I_{p}\left(\left(e_{y}-e_{x}\right)\circ f\right)\right|.$
Next, fix $s>1/\beta$ and $\eta\in(0,\beta-1/s)$. The Garsia-Rodemich-Rumsey inequality (see [GRR70] or [Hu17, Ch. 2]) applied with the Young function $u\mapsto u^{s}$ and exponent $\beta-\eta$ yields for any $g\in\mathcal{C}([0,T])$ that
$\sup_{\left|y-x\right\rvert\leq\varepsilon}\left|g(y)-g(x)\right\rvert\leq 8\varepsilon^{\beta-\eta-\frac{1}{s}}\left(\int_{0}^{T}\int_{0}^{T}\frac{\left|g(y)-g(x)\right\rvert^{s}}{\left|y-x\right\rvert^{1+s(\beta-\eta)}}\,dx\,dy\right)^{\frac{1}{s}}.$
By (32) with $s>1/\beta$ and the Kolmogorov continuity criterion, $I_{p}(f)$ admits a continuous modification, to which we can apply the Garsia-Rodemich-Rumsey inequality to get
(34) $\mathbb{E}\left(\omega(I_{p}(f),\varepsilon)\right)\leq 8\varepsilon^{\beta-\eta-\frac{1}{s}}\left(\int_{0}^{T}\int_{0}^{T}\frac{\mathbb{E}\left(\left|I_{p}\left(\left(e_{y}-e_{x}\right)\circ f\right)\right\rvert^{s}\right)}{\left|y-x\right\rvert^{1+s(\beta-\eta)}}\,dx\,dy\right)^{\frac{1}{s}}.$
Using (32) again, we can write
$\mathbb{E}\left(\left|I_{p}\left(\left(e_{y}-e_{x}\right)\circ f\right)\right\rvert^{s}\right)\leq\left(\left(p-1\right)^{\frac{s}{2}}\sqrt{p!}L\right)^{s}\left|y-x\right|^{\beta s}.$
Hence,
$\left(\int_{0}^{T}\int_{0}^{T}\frac{\mathbb{E}\left(\left|I_{p}\left(\left(e_{y}-e_{x}\right)\circ f\right)\right\rvert^{s}\right)}{\left|y-x\right\rvert^{1+s(\beta-\eta)}}\,dx\,dy\right)^{\frac{1}{s}}\leq\left(p-1\right)^{\frac{s}{2}}\sqrt{p!}L\left(\int_{0}^{T}\int_{0}^{T}\left|y-x\right\rvert^{\eta s-1}dxdy\right)^{\frac{1}{s}}.$
Plugged into (34), together with the straightforward computation
$\int_{0}^{T}\int_{0}^{T}\left|y-x\right\rvert^{\eta s-1}dxdy=\frac{2T^{1+\eta s}}{\eta s},$
yields
$\mathbb{E}\left(\omega(I_{p}(f),\varepsilon)\right)\leq 8\varepsilon^{\beta-\eta-\frac{1}{s}}\left(p-1\right)^{\frac{s}{2}}\sqrt{p!}L\left(\frac{2}{\eta s}\right)^{\frac{1}{s}}T^{\frac{1+\eta s}{s}}\varepsilon^{\beta-\eta-\frac{1}{s}}$
which combined with (33) concludes the proof. ∎
The next result shows how to bound the bounded Wasserstein distance between the law of a multiple Wiener-Itô integral valued in $\mathcal{D}$ and a centered Gaussian law on $\mathcal{D}$.
Proposition 5.7.
For a real separable Hilbert space $\mathfrak{H}$ and the Banach space $E=\mathcal{C}\left([0,T]\right)$ with $T>0$, let $f\in\gamma^{p}(\mathfrak{H},E)$ and $Z$ be a centered Radon Gaussian random element on $E$. Assume that both, $X$ and $Z$ almost surely take their values in the space of $\beta$-Hölder continuous functions on $[0,T]$. Denote the covariance operators of $F$ and $Z$ by $R_{F}$ and $R_{Z}$, respectively. Then there exists a positive constant $C$ depending on $p$, $T$, $\beta$ and the Hölder constants of $F$ and $Z$, such that
$d_{BL}\left(X,Z\right)\leq C\left(\left\lVert R_{F}-R_{Z}\right\rVert_{\nu}^{\alpha}+\sum_{r=1}^{p-1}\left\lVert f\otimes_{r}f\right\rVert_{\gamma^{2(p-r)}(\mathfrak{H},\mathfrak{N}(E^{\ast},E))}^{\alpha}\right),$
where $\alpha=\frac{\beta}{3\beta+2}$.
Proof.
We start by observing that by the triangle inequality,
$\mathbb{E}\left(\left\lVert\Gamma_{F}^{\pi}-Q_{Z}\right\rVert_{\pi}\right)$ $\leq\mathbb{E}\left(\left\lVert\Gamma_{F}^{\pi}-Q_{F}\right\rVert_{\pi}\right)+\mathbb{E}\left(\left\lVert Q_{F}-Q_{Z}\right\rVert_{\pi}\right)=\left\lVert Q_{F}-Q_{Z}\right\rVert_{\pi}+\mathbb{E}\left(\left\lVert\Gamma_{F}^{\pi}-Q_{F}\right\rVert_{\pi}\right),$
so that for any $\nu\leq 1$,
$\mathbb{E}\left(\left\lVert\Gamma_{F}^{\pi}-Q_{Z}\right\rVert_{\pi}\right)^{\nu}\leq\left\lVert Q_{F}-Q_{Z}\right\rVert_{\pi}^{\nu}+\mathbb{E}\left(\left\lVert\Gamma_{F}^{\pi}-Q_{F}\right\rVert_{\pi}\right)^{\nu}.$
Combining Proposition 4.7 and Theorem 5.4 yields the desired conclusion. ∎
6. Rates for a Gaussian-subordinated Hermite model
In this section, we consider a Gaussian-subordinated Hermite model in a fixed Wiener chaos of order $p$ on $\mathcal{C}([0,T])$. Such models are natural in Breuer–Major–type limits and functional CLTs for nonlinear transforms of Gaussian signals, and they appear in time–series and high–frequency settings where observables are polynomial functionals of approximately independent Gaussian innovations as well as in statistical physics and chaotic expansions, where one tracks how fixed-chaos components approximate Gaussian limits.
For $T>0$, let $Z$ be standard Brownian motion on $\mathcal{C}([0,T])$. Consider the Hilbert-Schmidt operator $K\colon L^{2}([0,T])\to L^{2}([0,T])$ defined by
$(Kf)(t)=\int_{0}^{T}R(s,t)f(s)ds,$
where $R$ is the covariance kernel of $Z$, given by $R(s,t)=\min\left\{s,t\right\}$. A Karhunen-Loève system (see [Loè78, Ch. XI] ) is an orthonormal basis $\{\varphi_{m}\}_{m\geq 1}$ of $L^{2}([0,T])$ and positive real numbers $\{\lambda_{m}\}_{m\geq 1}$ solving, for any $m\geq 1$,
(35) $K\varphi_{m}=\lambda_{m}\varphi_{m},$
so that $R$ admits the Mercer expansion
$R(s,t)=\sum_{m=1}^{\infty}\lambda_{m}\varphi_{m}(s)\varphi_{m}(t)$
in $L^{2}([0,T]^{2})$, and the Karhunen-Loève expansion of $Z$ reads
$Z=\sum_{m=1}^{\infty}\sqrt{\lambda_{m}}\varphi_{m}\zeta_{m},$
where the series converges $L^{2}(\Omega;\mathcal{C}([0,T]))$ and almost surely in $\mathcal{C}([[0,T]])$.
Differentiating (35) twice yields
$\lambda\varphi^{\prime\prime}+\varphi=0,\qquad\varphi(0)=0,\ \ \varphi^{\prime}(T)=0,$
which leads to the solutions
$\varphi_{m}(t)=\sqrt{\tfrac{2}{T}}\sin\!\Big(\tfrac{(m-\tfrac{1}{2})\pi t}{T}\Big),\qquad\lambda_{m}=\frac{T^{2}}{(m-\tfrac{1}{2})^{2}\pi^{2}},\qquad m\geq 1.$
For $k,m\in\mathbb{N}$, let
$\xi_{k,m}=H_{p}\left(W\left(\mathfrak{h}_{k,m}\right)\right)=I_{p}\left(h_{k,m}^{\otimes p}\right)\in\mathcal{H}_{p}(\mathbb{R}),$
where $H_{p}$ is the $p$-th Hermite polynomial, $W$ is an isonormal Gaussian process over a real, separable Hilbert space and $\left\{\mathfrak{h}_{k,m}\colon k,m\in\mathbb{N}\right\}$ is a doubly-indexed orthonormal basis of $\mathfrak{H}$ and, for $n\in\mathbb{N}$, define $\mathcal{C}([0,T])$-valued random elements $F_{n}$ by
$F_{n}=\frac{1}{\sqrt{p!}\,n}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\sqrt{\lambda_{m}}\varphi_{m}\,\xi_{k,m}.$
Then we have $F_{n}=I_{p}(f_{n})\in\mathcal{H}_{p}\left(\mathcal{C}([0,T])\right)$ with
(36) $f_{n}=\frac{1}{\sqrt{p!}\,n}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\sqrt{\lambda_{m}}\varphi_{m}\,\mathfrak{h}_{k,m}^{\otimes p}\in\gamma^{p}\left(\mathfrak{H},\mathcal{C}\left([0,T]\right)\right).$
Indeed,
$\left\lVert f_{n}\right\rVert_{\gamma^{p}\left(\mathfrak{H},\mathcal{C}\left([0,T]\right)\right)}^{2}$ $=\frac{1}{\sqrt{p!}\,n}\mathbb{E}\left(\left\lVert\sum_{k=1}^{n}\sum_{m=1}^{\infty}\sqrt{\lambda_{m}}\varphi_{m}\prod_{i=1}^{p}\zeta_{k,m}^{(i)}\right\rVert_{\infty}^{2}\right)$
$\leq\frac{1}{\sqrt{p!}\,n}\mathbb{E}\left(\left(\sum_{k=1}^{n}\sum_{m=1}^{\infty}\left(\prod_{i=1}^{p}\left|\zeta_{k,m}^{(i)}\right|\right)\lambda_{m}\left\lVert\varphi_{m}\right\rVert_{\infty}\right)^{2}\right)$
$=\frac{1}{\sqrt{p!}\,n}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\left\lVert\varphi_{m}\right\rVert_{\infty}^{2}$
$\leq\sqrt{\frac{2}{Tp!}}\sum_{m=1}^{\infty}\lambda_{m}<\infty.$
Applying our method, we obtain the following bound in the bounded Lipschitz distance.
Proposition 6.1.
In the above setting, for every $\zeta>0$ there exists a positive constant $C_{\zeta}$ such that
$d_{BL}\left(F_{n},Z\right)\leq C_{\zeta}\,n^{-\frac{1}{14}+\zeta}.$
Proof.
Denote the point evaluation in $x\in[0,T]$ on $\mathcal{C}\left([0,T]\right)$ by $e_{x}$. From (36), we immediately obtain
$\left\lVert\left(e_{s}-e_{t}\right)f_{n}\right\rVert_{\mathfrak{H}^{\otimes p}}^{2}=\frac{1}{np!}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\left(\varphi_{m}(s)-\varphi_{m}(t)\right)^{2}=\frac{1}{p!}\sum_{m=1}^{\infty}\lambda_{m}\left(\varphi_{m}(s)-\varphi_{m}(t)\right)^{2}.$
By the Karhunon-Loève identity,
$\sum_{m=1}^{\infty}\lambda_{m}\left(\varphi_{m}(s)-\varphi_{m}(t)\right)^{2}=s+t-2\min\left\{s,t\right\}=\left|t-s\right|,$
so that
$\left\lVert\left(e_{s}-e_{t}\right)f_{n}\right\rVert_{\mathfrak{H}^{\otimes p}}=\frac{1}{\sqrt{p!}}\left|t-s\right|^{\frac{1}{2}}.$
By Proposition 5.5, there exists a positive constant $C$ such that
$\mathbb{E}\left(\left\lVert F_{n,\varepsilon}-F_{n}\right\rVert_{\infty}\right)+\mathbb{E}\left(\left\lVert Z_{\varepsilon}-Z\right\rVert_{\infty}\right)\leq C\,\varepsilon^{\frac{1}{2}-\zeta}$
for any $\zeta\in(0,\frac{1}{2})$. As, by orthogonality,
$\mathbb{E}\left(\left(e_{s}-e_{t}\right)F_{n}\right)=\frac{1}{np!}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\varphi_{m}(s)\varphi_{m}(t)\mathbb{E}\left(\xi_{k,m}^{2}\right)=\sum_{m=1}^{\infty}\lambda_{m}\varphi_{m}(s)\varphi_{m}(t)=\min\left\{s,t\right\},$
the covariance kernels and hence also the covariance operators of $F_{n}$ and $Z$ coincide.
Turning to the contraction norms, we obtain by orthonormality of the $\mathfrak{h}_{k,m}$ that
$f_{n}\otimes_{r}f_{n}=\frac{1}{p!\,n}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\,\mathfrak{h}_{k,m}^{\otimes 2(p-r)}\otimes\varphi_{m}\otimes\varphi_{m},$
where we interpret the above as a projective tensor (as $\mathcal{C}\left([0,T]\right)$ has the approximation property, its projective tensor product is isometrically isomorphic to the corresponding space of nuclear operators). Denoting by $W^{(j)}$, $j\in\mathbb{N}$ independent copies of the isonormal Gaussian process $W$, we obtain
$\left\lVert f_{n}\otimes_{r}f_{n}\right\rVert^{2}_{\gamma^{2(p-r)}\left(\mathfrak{H},\mathcal{C}\left([0,T]\right)\widehat{\otimes}_{\pi}\mathcal{C}\left([0,T]\right)\right)}$
$\qquad\qquad=\frac{1}{n^{2}(p!)^{2}}\mathbb{E}\left(\left\lVert\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\,\left(\prod_{i=1}^{2(p-r)}W^{(i)}\left(\mathfrak{h}_{k,m}\right)\right)\varphi_{m}\otimes\varphi_{m}\right\rVert_{\pi}^{2}\right)$
$\qquad\qquad\leq\frac{1}{n^{2}(p!)^{2}}\sum_{k=1}^{n}\sum_{m=1}^{\infty}\lambda_{m}\left\lVert\varphi_{m}\right\rVert_{\infty}^{2}$
$\qquad\qquad=\frac{1}{n(p!)^{2}}\sum_{m=1}^{\infty}\lambda_{m}\left\lVert\varphi_{m}\right\rVert_{\infty}^{2}.$
Using that
$\frac{T^{2}}{2}=\int_{0}^{T}\mathbb{E}\left(Z(t)^{2}\right)\,dt=\sum_{m=1}^{\infty}\lambda_{m}$
and
$\left\lVert\varphi_{m}\right\rVert_{\infty}=\sqrt{\frac{2}{T}},$
we obtain
$\left\lVert f_{n}\otimes_{r}f_{n}\right\rVert^{2}_{\gamma^{2(p-r)}\left(\mathfrak{H},\mathcal{C}\left([0,T]\right)\widehat{\otimes}_{\pi}\mathcal{C}\left([0,T]\right)\right)}\leq\frac{T}{n(p!)^{2}},$
uniformly in $r\in\left\{1,2,\dots,p-1\right\}$. An application of Proposition 5.7 concludes the proof. ∎
7. The Hilbert space case
Let us briefly treat the special case where the Banach space $E$ is a separable Hilbert space. The following probability metric will be relevant:
Definition 7.1.
For a Banach space $E$, denote by $\mathcal{F}_{\infty}(E)\subseteq\mathcal{C}^{2}(E)$ the set of all twice Fréchet differentiable functions $f\colon E\to\mathbb{R}$ such that
$\left\lVert f^{\prime}\right\rVert_{\infty}+\left\lVert f^{\prime\prime}\right\rVert_{\infty}\leq 1.$
On the space of Borel probability measures on $E$, the metric $\rho_{\infty}$ is defined as
$\rho_{\infty}(\mu,\nu)=\sup_{f\in\mathcal{F}_{\infty}(E)}\left|\int_{E}f(x)\mu(dx)-\int_{E}f(x)\nu(dx)\right|.$
For Borel random elements $X$ and $Y$, we define $\rho_{\infty}(X,Y)=\rho_{\infty}(P_{X},P_{Y})$, where $P_{X}$ and $P_{Y}$ denote the laws of $X$ and $Y$.
It is well known that the $\rho_{\infty}$ distance does not metrize weak convergence on Banach spaces (separable or not), but it does metrize weak convergence on a separable Hilbert space – see [GL80] for a statement and proof of this fact, and [BBCP25] for details on weak convergence metrization on Hilbert spaces.
As all test functions $f$ in the defining class of the $\rho_{\infty}$-metric satisfy in particular $\left\lVert f^{\prime\prime}\right\rVert_{\infty}\leq 1$, the following result is an immediate consequence of Proposition 3.10.
Corollary 7.2.
Let $E$ be a Banach space, $X\in\mathbb{D}^{1,2}(E)$ be centered and let $Z$ be an independent, $E$-valued, centered Radon Gaussian random element on $E$ with covariance tensor $Q_{Z}$ and covariance operator $R_{Z}$. Then
$\rho_{\infty}(X,Z)\leq\frac{1}{2}\mathbb{E}\left(\left\lVert\Gamma_{\pi}\left(X,-L^{-1}X\right)-Q_{Z}\right\rVert_{\pi}\right).$
If $E$ has the approximation property, one also has
$\rho_{\infty}(X,Z)\leq\frac{1}{2}\mathbb{E}\left(\left\lVert\Gamma(X,-L^{-1}X)-R_{Z}\right\rVert_{\nu}\right).$
Let now $K$ be a separable Hilbert space. Our starting point is [DZ25, Thm. 3.1], which, stated in the notation of this paper, reads as follows.
Theorem 7.3.
Let $X\in\mathbb{D}^{1,2}(K)$ and let $Z$ be a non-degenerate Gaussian random variable on $K$ with covariance operator $Q_{Z}$. Then,
$\rho_{\infty}(X,Z)\leq\frac{1}{2}\left\lVert\Gamma(X,-L^{-1}X)-Q_{Z}\right\rVert_{L^{1}\left(\Omega;\mathcal{S}_{1}(K)\right)},$
where $\mathcal{S}_{1}(K)$ denotes the space of trace-class operators on $K$.
Remark 7.4.
Note that this result is also an immediate consequence of [BC20, Proof of Thm. 3.2], by applying the non-commutative $L^{1}-L^{\infty}$ Hólder inequality instead of the Cauchy-Schwarz inequality to the identity (3.5).
By recalling that in the Hilbert setting, $K\widehat{\otimes}_{\pi}K\simeq\mathcal{N}(K)\simeq\mathcal{S}_{1}(K)$ and $\gamma^{p}(\mathfrak{H},K)\simeq\mathfrak{H}^{\widehat{\odot}_{\sigma}p}\widehat{\otimes}_{\sigma}K$, it is immediate to observe that this theorem is a particular case of Theorem 7.2 (setting $E=K$). By making use of our contraction bound (Theorem 5.4), we thus obtain the following result.
Theorem 7.5.
For $p\in\mathbb{N}$ and $f\in\mathfrak{H}^{\widehat{\odot}_{\sigma}p}\widehat{\otimes}_{\sigma}K$, let $X=I_{p}(f)$. Furthermore, let $Z$ be a centered Gaussian random variable on $K$. Then
$\rho_{\infty}(F,Z)$ $\leq\frac{1}{2}\left\lVert R_{F}-R_{Z}\right\rVert_{\mathcal{S}_{1}(K)}+\frac{1}{2p}\sum_{r=1}^{p-1}\sqrt{(2p-2r)!}\alpha_{p,p,r}\left\lVert f\otimes_{r}f\right\rVert_{\gamma\left(\mathfrak{H}^{\widehat{\odot}_{\sigma}p},\mathcal{S}_{1}(K)\right)},$
where the constants $a_{p,p,r}$ are defined as in Theorem 5.4.